### v2

# InternVL3.5: Advancing Open-Source Multimodal Models in Versatility, Reasoning, and Efficiency

Abstract
[…]
* equal contribution; 🖂  corresponding authors; $\dagger$ technical leaders.
Figure 1: Comparison between InternVL3.5 and leading MLLMs in general capabilities. Hatched bars represent closed-source commercial models. We report average scores on a set of multimodal general, reasoning, text, and agentic benchmarks: MMBench v1.1 (en) liu2023mmbench , MMStar chen2024mmstar , BLINK fu2024blink , HallusionBench guan2023hallusionbench , AI2D kembhavi2016ai2d , OCRBench liu2023ocrbench , MMVet yu2023mmvet , MME-RealWorld (en) zhang2024mme , MVBench li2024mvbench , VideoMME fu2024video , MMMU yue2023mmmu , MathVista lu2023mathvista , MathVision wang2024mathvision , MathVerse zhang2024mathverse , DynaMath zou2024dynamath , WeMath qiao2024wemath , LogicVista xiao2024logicvista , MATH500 DBLP:conf/nips/HendrycksBKABTS21 , AIME24 aime2024 , AIME25 aime2025 , GPQA rein2024gpqa , MMLU-Pro wang2024mmlu , GAOKAO Zhang2023gaokao , IFEval zhou2023instruction , SGP-Bench qiu2024can , VSI-Bench yang2024think , ERQA erqa , SpaCE-10 gong2025space10 , and OmniSpatial jia2025omnispatial .
1 Introduction
The recent trend of Multimodal Large Language Models (MLLMs) qwen2.5 ; team2025keye_vl ; hong2025glm_v_thinking has gone beyond simple multimodal understanding and gradually focused on more general, complex, and realistic tasks such as text-related tasks aime2024 ; chen2021evaluating ; hendrycks2020measuring ; Zhang2023gaokao ; huang2023ceval ; rein2024gpqa , reasoning tasks yang2024think ; wang2024mathvision ; lu2023mathvista ; yue2023mmmu ; he2024olympiadbench and agentic tasks qin2025ui ; hu2025os_agents ; xie2024osworld ; zhang2025guimid ; yang2025zerogui ; sun2025scienceboard ; jia2025omnispatial ; yang2024think ; gong2025space10 . In these aspects, commercial models have created huge gaps with current open-source models, as shown in Table 2. Thereby, recent open-source efforts meng2025mm_eureka ; hong2025glm_v_thinking ; team2025keye_vl ; wang2025skywork_r1v2 aim to explore advanced reinforcement learning (RL) methods to mitigate the gap and pursue higher multimodal intelligence. However, despite much effort in RL algorithms shao2024deepseekmath ; zheng2025gspo ; meng2025mm_eureka ; cui2025prime ; cui2025entropy_mechanism ; yue2025does ; vl-rethinker ; xi2025bmmr and verifiers wang2025visualprm ; wang2023mathshepherd ; luo2024omegaprm , a stable, effective, and scalable reinforcement learning framework for MLLMs still remains an open problem in the community. Furthermore, the growth of multimodal capabilities, e.g., long visual context and high-resolution understanding zhu2025internvl3 ; llava_uhd_v2 ; luo2024llava_hr ; wang2025parameter ; bai2025qwen2_5 ; wang2025lvbenchextremelongvideo , often comes with ever increasing computational costs, which have become a crucial bottleneck of real-world applications.
In this work, we introduce InternVL3.5, an advanced family of InternVL series chen2023internvl ; chen2024internvl_1_5 ; chen2024internvl_2_5 ; zhu2025internvl3 ; luo2024mono_internvl ; mono_internvl_v1.5 ; gao2024mini_internvl with stronger capabilities in versatility, reasoning, and efficiency. Compared to InternVL3 zhu2025internvl3 , InternVL3.5 achieves superior performance through our proposed Cascade RL framework, which enhances reasoning capabilities in an efficient, scalable, and stable manner. Cascade RL consists of two complementary substages: an offline RL stage wang2024mpo , which efficiently achieves satisfactory performance, and an online RL stage zheng2025gspo , which carefully refines the output distribution and further push the performance upper bound of the model. The offline stage serves as an effective warm-up, ensuring high-quality rollouts for the subsequent online stage, thereby enabling the progressive improvement of MLLM reasoning abilities. In practice, Cascade RL demonstrates promising salability and stability, with a clear gain seen from InternVL3.5-1B to InternVL3.5-241B (Figure 5).
In addition, we equip InternVL3.5 with a much faster inference speed than its predecessor through two novel methods, namely Visual Resolution Router (ViR) and Decoupled Vision-Language Deployment (DvD). In particular, ViR aims to dynamically select the best trade-off resolution of visual tokens, thus reducing the inference costs with a negligible performance sacrifice. In practice, ViR can be efficiently integrated into InternVL3.5 with a light training stage namely Visual Consistency Learning (ViCO). Furthermore, DvD aims to deploy ViTs and LLMs on separate GPUs to maximize computational parallelism and hardware utilization. These two methods can be seamlessly combined to realize the hardware-friendly implementation for InternVL3.5.
[…]
In summary, our contributions include three folds:
(1) We release InternVL3.5, the latest family of the InternVL series with advanced reasoning abilities, powerful versatility, and promising efficiency. InternVL3.5 comprises various model scales (from 1B to 241B) with both dense and mixture-of-experts (MoE) models. All of our models and codes are publicly released.
(2) We propose three innovative designs in InternVL3.5, including cascade reinforcement learning (Cascade RL), visual resolution router (ViR) and decoupled vision-language deployment (DvD). These technologies significantly improve the capabilities and efficiency of InternVL3.5, providing practical tips to the community.
(3) We conduct extensive experiments and demonstrate that InternVL3.5 exhibits leading performance among open-source MLLMs hong2025glm_v_thinking ; team2025keye_vl ; coreteam2025mimovltechnicalreport ; yao2024minicpm ; wang2024qwen2vl . Compared to the latest commercial model, i.e., GPT-5 gpt5 , InternVL3.5 even achieves slightly better results on general multimodal capabilities. We believe our approach and open source will further advance the community.
2 InternVL3.5
Compared to its predecessors, the InternVL3.5 series achieves superior performance and faster inference. In Section 2.1, we introduce the model architectures of InternVL3.5 and InternVL3.5-Flash. For InternVL3.5-Flash, we further incorporate a Visual Resolution Router (ViR) module that dynamically selects the minimal resolution of visual tokens, achieving better inference efficiency. Section 2.2 and Section 2.3 describe the pre-training and post-training procedures of InternVL3.5, respectively. The details of our proposed Cascade Reinforcement Learning (Cascade RL) and Visual Consistency Learning (ViCO) methods are elaborated in Section 2.3. In Section 2.4, we present the test-time scaling approach used to further improve model performance. Finally, in Section 2.5, we describe the training and inference infrastructure supporting InternVL3.5, including implementation details of the Decoupled Vision-Language Deployment (DvD) framework. The overall architecture is shown in Figure 2, and the training recipes are illustrated in Figure 3.
Figure 2: Overall architecture. InternVL3.5 adopts the “ViT–MLP–LLM” paradigm as in previous versions. Building upon InternVL3.5, we further introduce InternVL3.5-Flash, which is extended with an additional visual resolution router (ViR) to dynamically select the appropriate compression rate (e.g., $\frac{1}{4}$ or $\frac{1}{16}$) for each image patch. Unlike Dynamic High Resolution which only splits image patches from the perspective of image width and height, our proposed ViR further introduces adaptivity from the perspective of semantic content.
Model Vision Encoder Language Model #Param
Vision Language Total
Dense Models
InternVL3.5-1B InternViT-300M Qwen3-0.6B 0.3B 0.8B 1.1B
InternVL3.5-2B InternViT-300M Qwen3-1.7B 0.3B 2.0B 2.3B
InternVL3.5-4B InternViT-300M Qwen3-4B 0.3B 4.4B 4.7B
InternVL3.5-8B InternViT-300M Qwen3-8B 0.3B 8.2B 8.5B
InternVL3.5-14B InternViT-300M Qwen3-14B 0.3B 14.8B 15.1B
InternVL3.5-38B InternViT-6B Qwen3-32B 5.5B 32.8B 38.4B
MoE Models
InternVL3.5-20B-A4B InternViT-300M GPT-OSS-20B 0.3B 20.9B 21.2B   (A4B)
InternVL3.5-30B-A3B InternViT-300M Qwen3-30B-A3B 0.3B 30.5B 30.8B   (A3B)
InternVL3.5-241B-A28B InternViT-6B Qwen3-235B-A22B 5.5B 235.1B 240.7B (A28B)
Table 1: Pre-trained models used in the InternVL3.5 series. “A” denotes the number of activated parameters.
2.1 Model Architecture
InternVL3.5. We follow the “ViT–MLP–LLM” paradigm adopted in previous versions of InternVL chen2024internvl_1_5 ; wang2024mpo ; chen2024internvl_2_5 ; zhu2025internvl3 ; gao2024mini_internvl . As shown in Table 1, we initialize the language model using the Qwen3 series yang2025qwen3 and GPT-OSS gpt-oss , and the vision encoder using InternViT-300M and InternViT-6B chen2023internvl . The Dynamic High Resolution strategy introduced in InternVL1.5 chen2024internvl_1_5 is also retained in our design.
InternVL3.5-Flash. Compared to InternVL3.5, InternVL3.5-Flash further integrates the Visual Resolution Router (ViR), thus yielding a series of efficient variants suitable for resource-constrained scenarios. Specifically, in InternVL3.5, each image patch is initially represented as 1024 visual tokens for the vision encoder, which are then compressed into 256 tokens via a pixel shuffle module before being passed to the Large Language Model (LLM). In InternVL3.5-Flash, as shown in Figure 2, an additional pixel shuffle module with a higher compression rate is included, enabling compression of visual tokens down to 64 tokens. For each patch, the patch router determines the appropriate compression rate by assessing its semantic richness, and routes it to the corresponding pixel shuffle module accordingly. Benefiting from this patch-aware compression mechanism, InternVL3.5-Flash is able to reduce the number of visual tokens by 50% while maintaining nearly 100% of the performance of InternVL3.5, as shown in Section 3.15.
2.2 Pre-Training
Training Objective. During the pre-training stage, we update all model parameters jointly using the combination of large-scale text and multimodal corpora. Specifically, given an arbitrary training sample consisting of a multimodal token sequence $\mathbf{x}=\left(x_{1},x_{2},\ldots,x_{L}\right)$, the next token prediction (NTP) loss radford2019language is calculated on each text token as follows:
$\mathcal{L}_{i}=-\log p_{\theta}\left(x_{i}\mid x_{1},\ldots,x_{i-1}\right),$ (1)
where $x_{i}$ is the predicted token and prefix tokens in $\{x_{1},x_{2},\ldots,x_{i-1}\}$ can be either text tokens or image tokens. In particular, for conversation samples, only response tokens are included for the loss calculation. Additionally, to mitigate bias toward either longer or shorter responses during training, we adopt the square averaging chen2024internvl_2_5 to reweight the NTP loss as follows:
$\mathcal{L}_{i}^{{}^{\prime}}=\frac{w_{i}}{\sum_{j}w_{j}}\cdot\mathcal{L}_{i},\quad w_{i}=\frac{1}{N^{0.5}},$ (2)
where $N$ denotes the number of tokens in the training sample on which the loss needs to be calculated. The random JPEG compression chen2024internvl_2_5 is also included to enhance the model’s real-world performance.
Data. The pre-training corpora can be classified into two categories: (1) Multimodal data: this subset of data is mainly sourced from the training corpora of InternVL3 zhu2025internvl3 , covering a diverse range of domains such as image captioning, general question answering, mathematics, scientific disciplines, charts, optical character recognition (OCR), knowledge grounding, document understanding, multi-turn dialogue, and medical data. (2) Text-only data: this part of data is constructed based on the training corpora of InternLM series 2023internlm ; cai2024internlm2 and is further augmented with open-source datasets benallal2024smollmcorpus ; acemath2024 ; scp116k ; muennighoff2025s1 . The pre-training corpora contains approximately 116M samples, corresponding to about 250B tokens. The ratio between text-only and multimodal data is approximately $1:2.5$. The maximum sequence length is set to 32K tokens to adapt long-context understanding and reasoning.
Figure 3: Training recipes of InternVL3.5. InternVL3.5 consists of three training stages: (1) native pre-training for vision-language alignment, (2) supervised fine-tuning for adaptation to downstream tasks, (3) Cascade RL for improvement on reasoning capabilities. InternVL3.5-Flash is an efficient version of InternVL3.5, which further integrates a visual resolution router (ViR) through consistency training and router training.
2.3 Post-Training
After the pre-training stage, we adopt a three-stage post-training strategy comprising: (1) Supervised Fine-Tuning (SFT), which maintains the same training objective as pre-training but leverages higher-quality conversation data to further enhance the model’s capabilities. (2) Cascade Reinforcement Learning (Cascade RL), which combines the benefits of offline and online RL methods to facilitate the reasoning capabilities. (3) Visual Consistency Learning (ViCO), which aims to integrate visual resolution router (ViR) into InternVL3.5 to construct InternVL3.5-Flash, by minimizing the output divergence of different visual compression rates.
Supervised Fine-Tuning. During the SFT phase, we adopt the same objective as in the pre-training stage and use the square averaging strategy chen2024internvl_2_5 to calculate the final loss. In this stage, the context window is set to 32K tokens to adapt long-context information. Compared to InternVL3, the SFT stage of InternVL3.5 contains more high-quality and diverse training data derived from three sources: (1) Instruction-following data from InternVL3, which are reused to preserve broad coverage of vision–language tasks. (2) Multimodal reasoning data in the “Thinking” mode, which are included to instill long-thinking capabilities in the model. To construct such data, we leverage a large-scale reasoning model to sample rollouts with detailed reasoning processes. In addition to validating whether answers are factually correct, we implement strict filtering measures for the reasoning processes themselves: this includes evaluating how clear the thinking is, weeding out redundancy, and ensuring that formatting remains consistent. The questions in these datasets cover various expert domains, such as mathematics and scientific disciplines, thereby strengthening performance on different reasoning tasks. (3) Capability-expansion datasets, which endow InternVL3.5 with new skills, including GUI-based interaction, embodied interaction, and scalable vector graphics (SVG) understanding and generation.
Cascade Reinforcement Learning. Compared to Pre-training and Supervised Fine-tuning (SFT), the core advantage of RL lies in its ability to introduce negative samples, which prune low-quality regions in the model’s output space and thereby enhance the overall response quality. As a derivative of the PPO algorithm schulman2017ppo , DPO rafailov2024dpo enables training based on existing rollouts, which we also regard as a form of offline RL. Offline RL algorithms rafailov2024dpo ; wang2024mpo often offer a higher training efficiency, but their performance ceiling is generally lower compared to online RL methods. In contrast, despite the effectiveness of online RL algorithms shao2024deepseekmath ; schulman2017ppo ; zheng2023ppo_max ; yu2025dapo , they are often computationally expensive and time-consuming. In this work, we propose Cascade RL, which aims to combine the benefits of offline RL and online RL to progressively facilitate the post-training of MLLMs in an efficient manner. Specifically, we first fine-tune the model using an offline RL algorithm wang2024mpo as an efficient warm-up stage to reach satisfied results, which can guarantee high-quality rollouts for the latter stage. Subsequently, we employ an online RL algorithm zheng2025gspo to further refine the output distribution based on rollouts generated by the model itself. Compared to the single offline or online RL stage, our cascaded RL achieves significant performance improvements at a fraction of the GPU time cost.
During the offline RL stage, we employ mixed preference optimization (MPO) wang2024mpo to fine-tune the model. Specifically, the training objective of MPO is a combination of preference loss $\mathcal{L}_{p}$, quality loss $\mathcal{L}_{q}$, and generation loss $\mathcal{L}_{g}$, which can be formulated as follows:
$\mathcal{L}_{\text{MPO}}=w_{p}\mathcal{L}_{p}+w_{q}\mathcal{L}_{q}+w_{g}\mathcal{L}_{g},$ (3)
where $w_{*}$ represents the weight assigned to each loss component. The DPO loss rafailov2024dpo , BCO loss jung2024bco , and LM loss brown2020gpt serve as the preference loss, quality loss, and generation loss, respectively.
During the online RL stage, we employ GSPO zheng2025gspo , without reference model constraints, as our online RL algorithm, which we find more effective in training both dense and mixture-of-experts (MoE) models. Similar to GRPO shao2024deepseekmath , the advantage is defined as the normalized reward across responses sampled from the same query:
$\widehat{A}_{i}=\frac{r\left(x,y_{i}\right)-\operatorname{mean}\left(\left\{r\left(x,y_{i}\right)\right\}_{i=1}^{G}\right)}{\operatorname{std}\left(\left\{r\left(x,y_{i}\right)\right\}_{i=1}^{G}\right)},$ (4)
where $y_{i}$ is the $i$-th response generated for the query $x$, $G$ is the total number of generated responses to the query, and $r\left(x,y_{i}\right)$ denotes the reward for this response. The training objective of GSPO is given by:
$\mathcal{L}_{\mathrm{GSPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\left\{y_{i}\right\}_{i=1}^{G}\sim\pi_{\theta\text{ old }}(\cdot\mid x)}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(s_{i}(\theta)\widehat{A}_{i},\operatorname{clip}\left(s_{i}(\theta),1-\varepsilon,1+\varepsilon\right)\widehat{A}_{i}\right)\right],$ (5)
where the importance sampling ratio is defined as the geometric mean of the per-token ratios:
$s_{i}(\theta)=\left(\frac{\pi_{\theta}\left(y_{i}\mid x\right)}{\pi_{\theta_{\text{old }}}\left(y_{i}\mid x\right)}\right)^{\frac{1}{\left|y_{i}\right|}}=\exp\left(\frac{1}{\left|y_{i}\right|}\sum_{t=1}^{\left|y_{i}\right|}\log\frac{\pi_{\theta}\left(y_{i,t}\mid x,y_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(y_{i,t}\mid x,y_{i,<t}\right)}\right),$ (6)
where $\pi_{\theta}\left(y_{i}\mid x,y_{i,<t}\right)$ and $\pi_{\theta}\left(y_{i,t}\mid x,y_{i,<t}\right)$ denote the generation probability of response $y_{i}$ and token $y_{i,t}$ under the policy model with parameters $\theta$, respectively.
Compared to directly training the model with a single RL paradigm, Cascade RL offers the following advantages: (1) Better training stability: In the offline RL stage, the rollout collection and parameter updates are decoupled, effectively mitigating issues such as reward hacking. During the online RL stage, we empirically observe that stronger models exhibit more stable and robust training dynamics. As a result, the performance gains achieved in the MPO stage further enhance the stability of the GSPO stage and reduce sensitivity to the algorithm. (2) Improved training efficiency: In the MPO stage, rollouts can be shared across different models, amortizing the sampling cost typically incurred during online RL. (3) Higher performance ceiling: Moreover, as shown in Section 3.15, models fine-tuned with MPO take fewer training steps to achieve higher performance in the subsequent online RL phase, further reducing training overhead.
Visual Consistency Learning. We further include ViCO as an additional training stage to integrate the visual resolution router (ViR) into InternVL3.5, thereby reducing the inference cost of InternVL3.5. The obtained efficient version of InternVL3.5 are termed as InternVL3.5-Flash. In particular, ViCO comprises two stages:
(1) Consistency training: In this stage, the entire model is trained to minimize the divergence between response distributions conditioned on visual tokens with different compression rates. In practice, we introduce an extra reference model, which is frozen and initialized with InternVL3.5. Given a sample, each image patch is represented as either 256 or 64 tokens, and the training objective is defined as follows:
$\mathcal{L}_{\text{ViCO}}=\mathbb{E}_{\xi\sim\mathcal{R}}\Bigg{[}\frac{1}{N}\sum_{i=1}^{N}\mathrm{KL}\Big{(}\pi_{\theta_{ref}}\left(y_{i}\mid y_{<i},I\right)\;\Big{\|}\;\pi_{\theta_{policy}}\left(y_{i}\mid y_{<i},I_{\xi}\right)\Big{)}\Bigg{]},$ (7)
where $\mathrm{KL}$ denotes the KL divergence and $\xi$ denotes the compression rate, which is uniformly sampled from $\{\frac{1}{4},\frac{1}{16}\}$. The image $I_{\xi}$ is represented as 256 tokens when $\xi=\frac{1}{4}$ and 64 tokens when $\xi=\frac{1}{16}$. We note that the reference model always performs inference with $\xi=\frac{1}{4}$.
(2) Router training: This stage aims to train the ViR to select an appropriate trade-off resolution for different inputs. ViR is formulated as a binary classifier and trained using standard cross-entropy loss. To construct the route targets, we first compute the KL divergence between the model outputs conditioned on uncompressed visual tokens (i.e., 256 tokens per patch) and those conditioned on compressed visual tokens (i.e., 64 tokens per patch). During this stage, the main MLLM (ViT, MLP and LLM) is kept frozen, and only the ViR is trained. Specifically, we first compute the loss ratio for each patch:
$r_{i}=\frac{\mathcal{L}_{\text{ViCO}}\big{(}y_{i}\mid I_{\frac{1}{16}}\big{)}}{\mathcal{L}_{\text{ViCO}}\big{(}y_{i}\mid I_{\frac{1}{4}}\big{)}},$ (8)
which quantifies the relative increase in loss caused by compressing the visual tokens. Based on this ratio, the binary ground-truth label for the patch router is defined as:
$y_{i}^{\text{router}}=\begin{cases}0,&r_{i}<\tau\;\text{(compression has negligible impact)}\\ 1,&r_{i}\geq\tau\;\text{(compression has significant impact)},\end{cases}$ (9)
where $y_{i}^{\text{router}}=0$ and $y_{i}^{\text{router}}=1$ indicate that the compression rate $\xi$ is set to $\tfrac{1}{16}$ and $\tfrac{1}{4}$, respectively. During training, we store the historical $r_{i}$ values of a sliding window, and $\tau$ is a dynamical threshold computed from the k-th percentile of historical $r_{i}$ values. In practice, the target distribution is balanced. During the consistency training stage, all patches of the same image are represented with a random compression rate, in order to ensure that the model retains its capability when no compression is applied. As shown in Section 3.15, InternVL3.5-Flash reduces 50% of the visual tokens while maintaining nearly 100% of the original performance.
Data. For the supervised fine-tuning (SFT) stage, the datasets comprise approximately 56 million samples, which corresponds to around 130 billion tokens. The proportion of text-only data to multimodal data is roughly 1:3.5. For the cascade reinforcement learning stage, we use MMPR-v1.2 wang2024mpo as the training data for offline RL, which contains about 200K sample pairs. Based on MMPR-v1.2, we compute the accuracy of each query using the provided rollouts and select those whose model accuracy falls between 0.2 and 0.8 for online RL. We further extend the dataset with recent multimodal datasets meng2025mm_eureka ; yang2025r1_onevision ; wang2025mv_math ; liu2024cmmmath ; deepscaler2025 to enhance diversity. The resulting dataset, termed MMPR-Tiny, consists of approximately 70K queries. We directly reuse the rollouts from MMPR-v1.2 for both offline RL and data filtering in online RL, thereby reducing the cost of sampling additional rollouts.
For the ViCO stage, we primarily leverage datasets identical to the SFT stage during consistency training, ensuring that the model retains its original performance. During router training, we use a subset of the SFT data, primarily composed of OCR and VQA examples, which are rich in visual information and sometimes require high-resolution understanding. This enables the resolution router to learn how to dynamically decide whether each image patch can be compressed based on the visual information.
2.4 Test-Time Scaling
Test-time scaling (TTS) has been empirically demonstrated as an effective approach to enhance the reasoning capabilities of LLMs and MLLMs, particularly for complex tasks that require multi-step inference wang2025visualprm ; snell2024test_time_scaling_efficient ; zhang2025stair ; luo2024omegaprm ; lightman2023prm800k . In this work, we implement a comprehensive test-time scaling approach that simultaneously improves reasoning depth (i.e., deep thinking) and breadth (i.e., parallel thinking). We note that unless otherwise specified, the experimental results reported in Section 3 are obtained without applying TTS. Thus far, we have only applied TTS to reasoning benchmarks, since we found that the model already exhibits strong perception and understanding capabilities, and initiating TTS yields no significant improvement.
Deep Thinking. By activating the Thinking mode, we guide the model to deliberately engage in step-by-step reasoning (i.e., decomposing complex problems into logical steps and validating intermediate conclusions) prior to generating the final answer. This approach systematically improves the logical structure of solutions for complex problems, particularly those requiring multi-step inference, and enhances reasoning depth.
Parallel Thinking. Following InternVL3, for reasoning tasks, we adopt the Best-of-N (BoN) strategy by employing VisualPRM-v1.1 wang2025visualprm as the critic model to select the optimal response from multiple reasoning candidates. This approach improves the reasoning breadth.
2.5 Infrastructure
Training Framework. The model training is conducted mainly based on the XTuner framework 2023xtuner , which incorporates a series of optimization strategies tailored for LLM and MoE training. These include fully shared data parallelism (FSDP) zhao2023fsdp to partition model parameters across GPUs, data packing chen2024internvl_2_5 to reduce padding tokens while balancing the token computation load across ranks for improved training efficiency, FP8 training based on DeepGEMM liu2024deepseekv3 and liger-kernel’s fused cross-entropy operator hsu2025ligerkernel to accelerate the training process, FlashAttention-3 dao2022flashattention ; dao2023flashattention2 to support packed inputs and speed up attention computation, and the TMA-Adaptive FP8 Grouped GEMM kernel suzhongling2025tmaadaptive to optimize the training of MoE models. For the online stage, we use verl sheng2024verl as our codebase. For InternVL3.5-20B-A4B, we implement an accelerated version of window attention with sink in GPT-OSS-20B through Triton tillet2019triton .
Figure 4: Overview of Decoupled Vision-Language Deployment. DvD decouples the vision and language models and deploys them on separate servers. The right side shows a time-consumption trace of the pipeline. (a) In the original deployment, the ViT, MLP, and LLM are executed sequentially. Given their substantial differences in size and computation patterns, this serial design significantly slows down inference. (b) With DvD, the inference of the ViT and the LLM is performed in parallel and asynchronously. Thus, ViT’s computations can overlap with the LLM’s prefilling and decoding, reducing resource conflicts and improving inference speed.
Decoupled Vision-Language Deployment. In multimodal inference, the vision encoder and language model have distinct computational characteristics. The vision encoder that transforms images into semantic features is highly parallelizable and does not rely on long-term history state. In contrast, the language model adopts the inference in an autoregressive manner, which requires previous states to compute the next one. This sequential property makes the language part more sensitive to memory bandwidth and latency. When MLLMs are deployed online at scale, the vision and language models often block each other, thus incurring additional inference cost. This effect becomes more pronounced with larger vision models or higher-resolution images.
As shown in Figure 4, we propose Decoupled Vision-Language Deployment (DvD) to address this issue by separating vision and language processing, with a particular focus on optimizing the prefilling stage. The vision subsystem batches and processes images to produce compact feature embeddings, which are then transmitted to the language subsystem for fusion with the text context prior to decoding. This separation alleviates blocking and brings multimodal prefilling performance closer to that of pure language models. In our system implementation, the ViT and MLP (and ViR for InternVL3.5-Flash) are deployed on the vision server, while the language server executes only the LLM. The communication is unidirectional, transmitting BF16 visual features over TCP, with RDMA optionally employed to achieve a higher transmission speed. Vision processing, feature transmission, and language processing are organized into an asynchronous three-stage pipeline, enabling overlapped execution and minimizing pipeline stalls.
DvD increases GPU utilization and processing efficiency on the vision side, while enabling the language server to focus exclusively on the LLM’s prefilling and decoding without being blocked by vision computation. This design leads to improved throughput and responsiveness. Moreover, the architecture supports independent hardware cost optimization for the vision and language modules, and facilitates the seamless integration of new modules without requiring modifications to the language server deployment.
3 Experiments
Here, we first compare the overall performance of InternVL3.5 with recent leading multimodal large language models (MLLMs) in Section 3.1. Subsequently, we evaluate our models in various domains, including multimodal reasoning and mathematics (Section 3.2), optical character recognition (OCR), chart and document understanding (Section 3.3), multi-image understanding (Section 3.4), real-world comprehension (Section 3.5), comprehensive multimodal evaluation (Section 3.6), multimodal hallucination evaluation (Section 3.7), visual grounding (Section 3.8), multimodal multilingual understanding (Section 3.9), video understanding (Section 3.10), GUI (Section 3.11), embodied (Section 3.12), and SVG (Section 3.13) tasks, most of which were tested using VLMEvalKit duan2024vlmevalkit . Additionally, the evaluation of the language capabilities of InternVL3.5 is presented in Section 3.14. Finally, we ablate newly proposed designs in InternVL3.5, including the Cascade Reinforcement Learning, Visual Resolution Router, and Decoupled Vision-Language Deployment (Section 3.15).
Task
Benchmark
InternVL3.5
InternVL3.5
GLM-4.1V
Kimi-VL-2506
Qwen2.5-VL
GLM-4.5V
Step-3
GPT-5
Size
30B-A3B
241B-A28B
9B
16B-A3B
72B
106B-A12B
321B-A38B
General
MMStar
72.0
77.9
72.9
70.4
70.8
75.3
69.0*
75.7*
MMVet
85.5
81.2
66.4*
78.1
76.2
75.2*
79.4*
77.6*
MMBench V1.1 (en)
84.8
87.4
85.8
84.4
88.4
88.2
81.1*
88.6*
MTVQA
33.7
39.3
25.5*
27.2*
31.7
30.5*
30.6*
33.1*
AI2D (w/ mask)
86.8
87.3
87.9
81.9
88.7
88.1
83.7*
89.5*
OCRBench
88.0
90.7
84.2
86.9
88.5
87.2
83.7*
80.7*
WildVision
75.8
82.8
74.0*
64.8*
78.6*
79.0*
89.4*
77.4*
MME-RealWorld (en)
64.8
65.1
61.7*
54.5*
63.2*
61.7*
54.0*
68.0*
HallusionBench
53.8
57.3
63.2
59.8
55.2
65.4
64.2
65.2*
MVBench
72.1
76.5
68.4
59.7
70.4
73.0
64.2*
74.0*
VideoMME (w/o sub)
68.7
72.9
68.2
67.8
73.3
74.6
63.6*
81.8*
MLVU
73.0
78.2
71.5*
74.2
74.6
75.3*
62.2*
77.3*
LongVideoBench
63.8
67.1
65.7*
64.5
60.7
68.8*
57.7*
72.6*
Overall
71.0
74.1
68.9
67.2
70.8
72.5
67.9
74.0
Reasoning
MMMU (val)
75.6
77.7
68.0
64.0
68.2
75.4
74.2
84.2
MathVista
80.9
82.7
80.7
80.1
74.2
84.6
79.2*
81.9*
MathVision
55.7
63.9
54.4
54.4
39.3
65.6
64.8
72.0*
MathVerse (vision-only)
60.4
68.5
68.4
54.6
47.3
72.1
62.7*
81.2*
DynaMath
36.5
46.5
42.5
28.1
35.9
53.9
50.1
60.9*
WeMath
48.4
62.3
63.8
42.0
49.1
68.8
59.8*
71.1*
OlympiadBench
62.9
68.7
56.3*
47.4
37.8*
64.0*
66.8*
73.2*
LogicVista
55.7
66.7
60.4
51.4
55.7
62.4
60.2*
70.0*
Overall
59.5
67.1
61.8
52.8
50.9
68.4
64.7
74.3
Text
MATH500
96.6
98.0
81.8*
91.8*
82.8*
94.2*
85.6*
97.8*
AIME24
79.4
84.7
36.2*
54.0*
15.0*
80.1*
86.6*
90.0*
AIME25
62.7
75.6
32.0*
39.1*
13.3*
72.8*
82.9
94.6
GPQA
68.2
73.2
50.3*
42.3*
52.0*
56.6*
73.0
85.7
MMLU-Pro
75.3
81.3
57.1*
68.5*
51.1*
69.7*
58.6*
85.6*
C-Eval
83.2
90.9
72.3*
64.4*
88.2*
89.1*
84.7*
88.2*
GAOKAO
91.9
94.5
78.4*
72.6*
92.9*
93.1*
70.2*
94.1*
IFEval
74.3
83.7
71.5*
65.8*
83.9*
82.4*
73.4*
94.6*
Overall
78.9
85.3
60.0
62.3
59.9
79.8
76.9
91.3
Agentic
SGP-Bench
69.4
70.7
57.1*
44.9*
57.1*
66.1*
56.5*
77.5*
ScreenSpot
86.6
89.8
87.1
ScreenSpot-v2
87.3
92.9
91.4
OSWorld-G
42.4
53.2
52.5
VSI-Bench
63.7
69.5
39.2*
37.4*
36.1*
41.4*
34.2*
37.5*
ERQA
41.5
46.8
45.8
36.0
44.8
46.5
44.5
65.7*
SpaCE-10
45.5
55.0
43.4*
39.2*
37.9*
51.6*
42.6*
43.8*
OmniSpatial
48.1
51.9
47.7
37.3
47.9
51.0
47.0
59.6*
Overall
60.6
66.2
Table 2: The overall comparison of InternVL3.5 series and existing open-source and closed-source MLLMs. *: reproduced through VLMEvalkit duan2024vlmevalkit . : reported by GLM-4.5V hong2025glm_v_thinking . : reported by OpenCompass opencompass2023 .
3.1 Overall Comparison with Other Advanced MLLMs
Table 2 presents a comprehensive evaluation of InternVL3.5’s performance across 35 benchmarks categorized into four key multimodal task types: (1) General Tasks: MMStar chen2024mmstar , MMVet yu2023mmvet , MMBench V1.1 (en) liu2023mmbench , MTVQA tang2024mtvqa , AI2D kembhavi2016ai2d , OCRBench liu2023ocrbench , WildVision lu2024wildvision , MME-RealWorld (en) zhang2024mme , HallusionBench guan2023hallusionbench , MVBench li2024mvbench , VideoMME fu2024video , MLVU MLVU , (2) Reasoning Tasks: MMMU yue2023mmmu , MathVista lu2023mathvista , MathVision lu2023mathvista , MathVerse zhang2024mathverse , DynaMath zou2024dynamath , WeMath qiao2024wemath , OlympiadBench he2024olympiadbench , LogicVista xiao2024logicvista ; (3) Text-Centric Tasks: MATH500 lightman2023prm800k , AIME24 aime2024 , AIME25 aime2025 , GPQA rein2024gpqa , MMLU-Pro li2023cmmlu , GAOKAO Zhang2023gaokao , IFEval zhou2023instruction ; (4) Agentic Tasks: SGP-Bench qiu2024can , ScreenSpot cheng2024seeclick , ScreenSpot-v2 wu2024atlas , OSWorld-G xie2024osworld , VSI-Bench yang2024think , ERQA erqa , SpaCE-10 gong2025space10 .
We report results of our flagship models (InternVL3.5-30B-A3B and InternVL3.5-241B-A28B) and frontier open-source MLLMs (GLM-4.1V hong2025glm_v_thinking , Kimi-VL-A3B-2506 team2025kimi , GLM-4.5V hong2025glm_v_thinking , Qwen2.5-VL-72B bai2025qwen2_5 and Step-3 wang2025step ). We also include a state-of-the-art closed-source MLLM (GPT-5 gpt5 ) for comparison.
These results highlight InternVL3.5’s strong capabilities across diverse tasks. For general multimodal tasks, InternVL3.5 demonstrates leading performance among open-source models on general multimodal understanding (e.g., MMVet), multilingual (e.g., MTVQA), OCR (e.g., OCRBench), real-world (e.g., MME-RealWorld), and video (e.g., LongVideoBench) benchmarks. It even achieves a similar overall score as GPT-5 gpt5 (74.1 vs. 74.0), the state-of-the-art closed-source MLLM. Nevertheless, some tasks like HallusionBench still pose challenges, indicating the need for further refinements.
We also observe particularly significant gains in complex multimodal reasoning, as evidenced by the scores of 77.7 on MMMU and 82.7 on MathVista, which surpass most open-source models and approaching top-tier commercial systems. These improvements are largely driven by enhanced training strategies, especially Cascade RL, and refined test-time scaling methodologies, enabling robust generalization in challenging domains such as mathematical reasoning (e.g., MathVerse) and multidisciplinary understanding (e.g., MTVQA).
For text-related tasks, InternVL3.5 outperforms most open-source models with significant margins. This is primarily due to our native pre-training strategy introduced in InternVL3 zhu2025internvl3 , where we include a large amount of text-only data in the training process. This strategy allows the model to simultaneously acquire linguistic and multimodal abilities in a more efficient and integrated manner, and preserve the language capabilities of the pre-trained LLM to avoid the catastrophic forgetting issue luo2024mono_internvl . As a result, InternVL3.5 achieves high performance on both general (e.g., IFEval) and reasoning (e.g., AIME24 and MMLU-Pro) benchmarks and narrows the gap with GPT-5.
We also demonstrate the versatility of InternVL3.5 through agentic benchmarks. On SGP-Bench, an SVG understanding benchmark, InternVL3.5 achieves leading performance (69.4 and 70.7) that surpasses all open-source models. For GUI tasks, our models show strong abilities on GUI grounding (ScreenSpot) and online agentic (OSWorld-G) benchmarks. Evaluation on embodied tasks (VSI-Bench, ERQA, SpaCE-10, OmniSpatial) validates InternVL3.5’s spatial reasoning skills and the competence to understand complex and dynamic environments, highlighting its potential in embodied agents, robotic navigation, and interactive scene perception.
3.2 Multimodal Reasoning and Mathematics
To comprehensively evaluate the multimodal reasoning and mathematical capabilities of InternVL3.5, we conduct extensive experiments across a series of benchmarks, including MMMU yue2023mmmu for multidisciplinary reasoning, MathVista lu2023mathvista , MathVision wang2024mathvision , and MathVerse zhang2024mathverse for mathematical reasoning, as well as DynaMath zou2024dynamath , WeMath qiao2024wemath , and LogicVista xiao2024logicvista for complementary logical reasoning assessment.
As shown in Table 3, InternVL3.5 achieves state-of-the-art performance across all evaluated benchmarks among open-source models. Compared with the previous generation InternVL3, InternVL3.5 improves reasoning performance by more than 10 points across all model sizes relative to their counterparts of the comparable scale. Furthermore, InternVL3.5-241B-A28B consistently outperforms all open-source counterparts, obtaining an overall average score of 66.9. It is followed by InternVL3.5-38B, which secures the second highest score with an average of 66.0. At mid-scale, the performance of InternVL3.5 is particularly impressive, with InternVL3.5-30B-A3B and InternVL3.5-14B attaining scores of 75.6 and 73.3 on MMMU respectively, both outperforming the larger InternVL3-78B (72.2). At lightweight scales, InternVL3.5 achieves substantial gains over open-source baselines. Compared to its predecessor InternVL3, it delivers marked improvements: the InternVL3.5-2B model has an average score of 50.7, significantly higher than that of InternVL3-2B with a score of 32.4; the InternVL3.5-4B model, with a score of 57.4, far exceeds the score of 33.5 of MiniCPM-V-4; and the InternVL3.5-8B model achieves a score of 60.3, substantially surpassing the score of 44.3 of InternVL3-8B. These significant improvements are mainly from our Cascade RL, showing its strong scalability for reasoning tasks. The ablation study about how different training stages influence the reasoning abilities of our models is presented in Section 3.15.
Furthermore, our experiments also demonstrate that Cascade RL can be seamlessly combined with parallel thinking and obtain further gains. For instance, with parallel thinking, the overall reasoning scores of InternVL3.5-4B, InternVL3.5-8B and InternVL3.5-241B-A28B are further improved by +2.6%, +2.1% and +1.8%, respectively, highlighting the effectiveness of test-time scaling for reasoning-related tasks.
Model MMMU (val) MathVista (mini) MathVision MathVerse (vision-only) DynaMath (worst case) WeMath LogicVista Overall
InternVL3-1B zhu2025internvl3 43.4 45.8 18.8 18.7 5.8 13.4 29.8 25.1
InternVL3.5-1B 44.2 59.3 27.3 37.8 17.2 21.5 29.3 33.8
w/ Parallel Thinking wang2025visualprm 51.0 69.5 33.4 45.8 25.0 36.4 41.8 43.3
Ovis-2B lu2024ovis 45.6 64.1 17.7 29.4 10.0 9.9 34.7 30.2
Qwen2.5-VL-3B bai2025qwen2_5 51.2 61.2 21.9 31.2 13.2 22.9 40.3 34.6
InternVL3-2B zhu2025internvl3 48.6 57.0 21.7 25.3 14.6 22.4 36.9 32.4
InternVL3.5-2B 59.0 71.8 42.8 53.4 31.5 48.5 47.7 50.7
w/ Parallel Thinking wang2025visualprm 64.8 74.3 49.0 54.2 33.3 52.6 49.4 53.9
Ovis-4B lu2024ovis 49.0 69.6 21.5 38.5 18.0 16.9 35.3 35.5
MiniCPM-V-4-4B yao2024minicpm 51.2 66.9 20.7 18.3 14.2 32.7 30.6 33.5
InternVL2.5-4B chen2024internvl_2_5 51.8 64.1 18.4 27.7 15.2 21.2 34.2 33.2
InternVL3.5-4B 66.6 77.1 54.4 61.7 35.7 50.1 56.4 57.4
w/ Parallel Thinking wang2025visualprm 71.4 79.2 57.5 60.0 38.3 53.0 60.4 60.0
MiniCPM-o2.6 yao2024minicpm 50.9 73.3 21.7 35.0 10.4 25.2 36.0 36.1
Ovis-8B lu2024ovis 57.4 71.8 25.9 42.3 20.4 27.2 39.4 40.6
Qwen2.5-VL-8B bai2025qwen2_5 55.0 67.8 25.4 41.1 21.0 35.2 44.1 41.4
MiMo-VL-RL-8B coreteam2025mimovltechnicalreport 66.7 81.5 60.4 71.5 45.9 66.3 61.4 64.8
Keye-VL-8B team2025keye_vl 71.4 80.7 50.8 54.8 37.3 60.7 54.8 58.6
GLM-4.1V-9B hong2025glm_v_thinking 68.0 80.7 54.4 68.4 42.5 63.8 60.4 62.6
InternVL3-8B zhu2025internvl3 62.7 71.6 29.3 39.8 25.5 37.1 44.1 44.3
InternVL3.5-8B 73.4 78.4 56.8 61.5 37.7 57.0 57.3 60.3
w/ Parallel Thinking wang2025visualprm 73.4 80.8 59.9 62.6 39.9 57.6 62.6 62.4
Gemma-3-12B team2025gemma 55.2 56.1 30.3 21.1 20.8 33.6 41.2 36.9
Ovis2-16B lu2024ovis 60.7 73.7 30.1 45.8 26.3 45.0 47.4 47.0
InternVL3-14B zhu2025internvl3 67.1 75.1 37.2 44.4 31.3 43.0 51.2 49.9
InternVL3.5-14B 73.3 80.5 59.9 62.8 38.7 58.7 60.2 62.0
w/ Parallel Thinking wang2025visualprm 74.3 81.5 61.4 64.1 41.3 60.8 61.3 63.5
Kimi-VL-A3B-2506 team2025kimi 64.0 80.1 54.4 54.6 28.1 42.0 51.4 53.5
InternVL3.5-20B-A4B 72.6 78.0 53.0 57.0 33.1 41.4 56.8 56.0
w/ Parallel Thinking wang2025visualprm 74.0 79.2 55.7 58.9 35.9 45.5 58.2 58.2
InternVL3.5-30B-A3B 75.6 80.9 55.7 60.4 36.5 48.4 55.7 59.0
w/ Parallel Thinking wang2025visualprm 75.0 82.1 58.5 61.4 38.3 57.5 59.7 61.8
Gemma-3-27B team2025gemma 64.9 59.8 39.8 34.0 28.5 37.9 47.3 44.6
Ovis2-34B lu2024ovis 66.7 76.1 31.9 50.1 27.5 51.9 49.9 50.6
Qwen2.5-VL-32B bai2025qwen2_5 70.2 74.8 38.1 57.6 35.1 46.5 52.6 53.6
Skywork-R1V3-38B shen2025skywork_r1v3 76.0 77.1 52.6 59.6 35.1 56.5 59.7 59.5
InternVL3-38B zhu2025internvl3 70.1 75.1 34.2 48.2 35.3 48.6 58.4 52.8
InternVL3.5-38B 76.9 81.9 63.7 67.6 41.7 64.8 65.3 66.0
w/ Parallel Thinking wang2025visualprm 76.7 83.8 65.6 66.5 44.9 67.8 64.7 67.1
GPT-5-nano-20250807 gpt5 72.6 73.1 59.7 66.6 47.9 59.4 57.5 62.4
GPT-5-20250807 gpt5 81.8 81.9 72.0 81.2 60.9 71.1 70.0 74.1
Claude-3.7-Sonnet claude3series2024 75.0 66.8 41.9 46.7 39.7 49.3 58.2 53.9
Gemini-2.0-Pro gemini2_0 69.9 71.3 48.1 67.3 43.3 56.5 53.2 58.5
Gemini-2.5-Pro gemini2_0 74.7 80.9 69.1 76.9 56.3 78.0 73.8 72.8
Doubao-1.5-Pro guo2025seedvl1_5 73.8 78.6 51.5 64.7 44.9 65.7 64.2 63.3
GLM-4.5V hong2025glm_v_thinking 75.4 84.6 65.6 72.1 53.9 68.8 62.4 69.0
QvQ-72B-Preview qvq-72b-preview 70.3 70.3 34.9 48.2 30.7 39.0 58.2 50.2
Qwen2.5-VL-72B bai2025qwen2_5 68.2 74.2 39.3 47.3 35.9 49.1 55.7 52.8
Step3-321B-A38B wang2025step 74.2 79.2 64.8 62.7 50.1 59.8 60.2 64.4
InternVL3-78B zhu2025internvl3 72.2 79.0 43.1 51.0 35.1 46.1 55.9 54.6
InternVL3.5-241B-A28B 77.7 82.7 63.9 68.5 46.5 62.3 66.7 66.9
w/ Parallel Thinking wang2025visualprm 78.7 84.8 65.9 71.6 47.8 64.4 67.6 68.7
Table 3: Comparison of multimodal reasoning and mathematical performance. MMMU yue2023mmmu is a multidisciplinary reasoning benchmark. MathVista lu2023mathvista , MathVision wang2024mathvision , MathVerse zhang2024mathverse , DynaMath zou2024dynamath , and WeMath qiao2024wemath are mathematics benchmarks. LogicVista xiao2024logicvista is a logical reasoning benchmark. Part of the results are collected from other papers team2025keye_vl ; hong2025glm_v_thinking ; coreteam2025mimovltechnicalreport ; wang2024qwen2vl ; zhu2025internvl3 and the OpenCompass leaderboard opencompass2023 . The overall score is the average score of all benchmarks.
3.3 OCR, Chart, and Document Understanding
To evaluate the comprehensive capabilities of the model across tasks related to text, document, and chart comprehension, we conduct an extensive assessment on nine benchmarks: AI2D kembhavi2016ai2d , ChartQA masry2022chartqa , TextVQA singh2019textvqa , DocVQA mathew2021docvqa , InfoVQA mathew2022infographicvqa , OCRBench liu2023ocrbench , SEED-2-Plus li2024seedbench2plus , CharXiv wang2024charxiv , and VCR zhang2024vcr .
As shown in Table 4, InternVL3.5 achieves competitive results on these benchmarks, outperforming other open-source and closed-source models. At the lightweight scale, InternVL3.5 demonstrates significant potential. For instance, InternVL3.5-2B attains an overall average score of 76.7 across nine benchmarks, surpassing InternVL3-2B of similar size, which scores 74.7. Specifically, on DocVQA, InfoVQA, and SEED-2-Plus, InternVL3.5-2B achieves scores of 89.4, 70.8, and 68.0, respectively, compared to 88.3, 66.1, and 64.6 for InternVL3-2B, highlighting InternVL3.5’s strong vision-language understanding capabilities.
Moreover, as the model scale increases, InternVL3.5 continues to deliver enhanced performance in vision-language understanding. Across the same nine benchmarks, InternVL3.5-4B, InternVL3.5-20B-A4B, InternVL3.5-14B, InternVL3.5-30B-A3B, and InternVL3.5-38B achieve overall average scores of 80.0, 81.7, 82.0, 83.9, and 84.6, respectively, demonstrating consistent improvements with larger model sizes. On AI2D, InternVL3.5-2B, InternVL3.5-4B, InternVL3.5-14B and InternVL3.5-38B obtain scores of 78.8/89.1, 82.6/92.3, 85.1/93.3, and 87.8/95.1, respectively, indicating a consistent upward trend across larger models. On ChartQA, InternVL3.5-4B shows a notable improvement in chart understanding, increasing from 80.7 (InternVL3.5-2B) to 86.0, reflecting a significant gain in visual reasoning capability. Similarly, on CharXiv and VCR-EN-Easy, larger model variants also achieve measurable improvements in text and document understanding.
Model AI2D (w / wo M) ChartQA (test avg) TextVQA (val) DocVQA (test) InfoVQA (test) OCR Bench SEED-2 Plus CharXiv (RQ / DQ) VCR-EN-Easy (EM / Jaccard) Overall
InternVL3-1B zhu2025internvl3 69.4 / 78.3 75.3 74.1 81.9 53.7 790 58.2 21.0 / 47.1 89.3 / 96.2 68.6
InternVL3.5-1B 71.1 / 81.8 77.7 71.5 85.6 60.5 795 62.3 26.9 / 60.6 83.5 / 94.0 71.3
Qwen2-VL-2B wang2024qwen2vl 74.7 / 84.6 73.5 79.7 90.1 65.5 809 62.4 81.5 /   –
Aquila-VL-2B gu2024aquilavl 75.0 /   – 76.5 76.4 85.0 58.3 772 63.0 70.0 /   –
Qwen2.5-VL-3B bai2025qwen2_5 81.6 /   – 84.0 79.3 93.9 77.1 797 67.6 31.3 / 58.6
InternVL3-2B zhu2025internvl3 78.7 / 87.4 80.2 77.0 88.3 66.1 835 64.6 28.3 / 54.7 91.2 / 96.9 74.7
InternVL3.5-2B 78.8 / 89.1 80.7 76.5 89.4 70.8 836 68.0 31.6 / 65.0 90.1 / 96.4 76.7
MiniCPM-V-4-4B yao2024minicpm 80.9 / 91.4 73.0 81.4 94.0 67.0 862 67.0 31.9 / 56.4 80.9 / 90.1 75.0
InternVL3.5-4B 82.6 / 92.3 86.0 77.9 92.4 78.0 822 69.4 39.6 / 71.1 91.6 / 97.0 80.0
Ovis1.6-Gemma2-9B lu2024ovis 84.4 /   – 830
MiniCPM-V2.6-8B yao2024minicpm 82.1 /   – 82.4 80.1 90.8 852 65.7 31.0 / 57.1 73.9 / 85.7
Molmo-7B-D deitke2024molmo   –    / 93.2 84.1 81.7 92.2 72.6 694
Qwen2-VL-7B wang2024qwen2vl 83.0 / 92.1 83.0 84.3 94.5 76.5 866 69.0 89.7 / 93.8
Qwen2.5-VL-7B bai2025qwen2_5 83.9 /   – 87.3 84.9 95.7 82.6 864 70.4 42.5 / 73.9
Keye-VL-8B team2025keye_vl 85.8 / 88.5 72.5 75.7 87.0 63.0 853 67.8 36.8 / 75.2
GLM-4.1V-9B hong2025glm_v_thinking 82.2 / 87.0 70.0 79.6 93.3 80.3 823 71.8 53.4 / 82.4 32.7 / 55.2 72.5
InternVL3-8B zhu2025internvl3 85.2 / 92.6 86.6 80.2 92.7 76.8 880 69.7 37.6 / 73.6 94.5 / 98.1 81.3
InternVL3-9B zhu2025internvl3 84.6 / 92.9 86.2 79.4 93.6 79.6 877 68.8 38.0 / 72.5 94.2 / 97.9 81.3
InternVL3.5-8B 84.0 / 92.8 86.7 78.2 92.3 79.1 840 70.8 44.4 / 72.2 92.6 / 97.3 81.2
Gemma3-12B team2025gemma 84.2 / 87.2 75.7 67.7 87.1 64.9 702 65.5 24.9 / 65.4
InternVL3-14B zhu2025internvl3 86.0 / 93.7 87.3 80.5 94.1 83.6 875 70.3 43.1 / 82.2 94.8 / 98.2 83.4
InternVL3.5-14B 85.1 / 93.3 86.5 77.8 93.4 78.3 836 70.7 47.9 / 76.6 93.4 / 97.7 82.0
Kimi-VL-A3B-2506 team2025kimi 81.9 / 91.2 73.7 77.7 93.5 74.5 869 70.8 46.8 / 71.5 81.1 / 90.6 78.3
InternVL3.5-20B-A4B 85.9 / 93.5 86.6 78.5 92.9 78.1 870 69.3 38.2 / 78.5 93.7 / 97.8 81.7
InternVL3.5-30B-A3B 86.8 / 94.5 87.4 80.5 94.2 81.4 880 70.6 48.0 / 81.8 94.9 / 98.2 83.9
Gemma3-27B team2025gemma 84.5 / 88.2 78.0 65.1 86.6 65.1 717 68.1 25.7 / 63.8
Qwen2.5-VL-32B bai2025qwen2_5 84.0 / 92.9 64.9 78.9 94.8 83.4 850 72.1 46.1 / 84.9 88.9 / 93.0 80.7
Cambrian-34B tong2024cambrian 79.5 /   – 75.6 76.7 75.5 46.0 600 27.3 / 59.7 79.7 / 89.3
VILA-1.5-40B lin2024vila 69.9 /   – 67.2 73.6 460 24.0 / 38.7
InternVL3-38B zhu2025internvl3 88.9 / 95.5 89.2 83.9 95.4 85.0 886 71.6 46.4 / 87.2 96.1 / 98.7 85.5
InternVL3.5-38B 87.8 / 95.1 88.8 82.7 94.0 83.8 870 71.0 48.1 / 83.1 95.3 / 98.4 84.6
GPT-4V gpt4v 78.2 / 89.4 78.5 78.0 88.4 75.1 645 53.8 37.1 / 79.9 52.0 / 65.4 70.0
GPT-4o-20240513 chatgpt4o 84.6 / 94.2 85.7 77.4 92.8 79.2 736 72.0 47.1 / 84.5 91.6 / 96.4 81.6
Claude-3-Opus claude3series2024 70.6 / 88.1 80.8 67.5 89.3 55.6 694 44.2 30.2 / 71.6 62.0 / 77.7 67.3
Claude-3.5-Sonnet claude3series2024 81.2 / 94.7 90.8 74.1 95.2 74.3 788 71.7 60.2 / 84.3 63.9 / 74.7 78.7
Gemini-1.5-Pro reid2024gemini1_5 79.1 / 94.4 87.2 78.8 93.1 81.0 754 43.3 / 72.0 62.7 / 77.7
GLM-4.5V hong2025glm_v_thinking 88.1 / 93.7 86.6 72.0 94.5 84.1 872 74.0 59.5 / 88.2 36.5 / 44.9 75.8
NVLM-D-72B dai2024nvlm 85.2 / 94.2 86.0 82.1 92.6 853
Molmo-72B deitke2024molmo   –    / 96.3 87.3 83.1 93.5 81.9
Qwen2-VL-72B wang2024qwen2vl 88.1 /   – 88.3 85.5 96.5 84.5 877 91.3 / 94.6
Qwen2.5-VL-72B bai2025qwen2_5 88.7 /   – 89.5 83.5 96.4 87.3 885 73.0 49.7 / 87.4
InternVL3-78B zhu2025internvl3 89.7 / 96.0 89.7 84.3 95.4 86.5 906 71.9 46.0 / 85.1 96.0 / 98.6 85.8
InternVL3.5-241B-A28B 87.3 / 95.0 88.0 84.5 94.9 82.0 907 71.9 48.3 / 83.9 97.0 / 99.0 85.2
Table 4: Comparison of OCR, chart, and document understanding performance. We compare OCR-related performance across 9 benchmarks: AI2D kembhavi2016ai2d , ChartQA masry2022chartqa , TextVQA singh2019textvqa , DocVQA mathew2021docvqa , InfoVQA mathew2022infographicvqa , OCRBench liu2023ocrbench , SEED-2-Plus li2024seedbench2plus , CharXiv wang2024charxiv , and VCR zhang2024vcr . Part of results are collected from the OpenCompass leaderboard opencompass2023 and other papers dubey2024llama3 ; deitke2024molmo ; claude3series2024 ; wang2024charxiv ; zhang2024vcr . When calculating Overall, the score of OCR Bench is normalized from 0-1000 to 0-100.
Model Name BLINK (val) Mantis Eval MMIU Muir Bench MMT (val) MIRB (avg) Overall RealWorld QA MME-RW (EN) WildVision (win rate) R-Bench (dis) Overall
InternVL3-1B zhu2025internvl3 42.9 50.2 39.3 31.2 52.9 36.1 42.1 58.2 46.0 43.8 60.4 52.1
InternVL3.5-1B 44.0 54.8 45.2 41.7 54.5 44.2 47.4 57.6 46.8 49.2 57.4 50.6
Qwen2-VL-2B wang2024qwen2vl 44.4 55.1 62.6
Qwen2.5-VL-3B bai2025qwen2_5 47.6 47.7 65.4 53.1
InternVL3-2B zhu2025internvl3 50.3 65.9 43.0 38.8 59.5 42.9 50.1 64.3 53.8 48.8 67.5 58.6
InternVL3.5-2B 51.3 58.5 44.9 44.0 58.5 45.9 50.5 62.0 49.7 66.6 62.4 58.5
MiniCPM-V-4 yao2024minicpm 53.0 71.4 46.1 59.7 66.0 58.4 46.0 64.2 58.7
InternVL3.5-4B 58.1 62.7 49.2 53.1 64.3 55.9 57.2 66.3 59.8 69.8 68.7 65.1
Qwen2-VL-7B wang2024qwen2vl 53.2 64.0 70.1 56.5 64.0
Qwen2.5-VL-7B bai2025qwen2_5 56.4 59.6 68.5 57.4
MiniCPM-V2.6 yao2024minicpm 53.0 69.0 60.8 65.0
Keye-VL-8B team2025keye_vl 53.4 50.4 65.4 65.9 48.8 66.6 65.7 61.7
GLM-4.1V-9B hong2025glm_v_thinking 65.9 74.7 68.4 70.6 61.7 74.0 69.1 68.9
InternVL3-8B zhu2025internvl3 55.5 70.1 46.8 55.0 65.0 56.8 58.2 70.8 62.0 69.8 74.1 69.2
InternVL3-9B zhu2025internvl3 58.6 70.1 50.4 51.4 65.4 58.6 59.1 70.5 61.3 63.8 70.3 66.5
InternVL3.5-8B 59.5 70.5 49.4 55.8 66.7 57.3 59.9 67.5 62.8 76.6 69.7 67.4
Gemma3-12B team2025gemma 52.6 49.7 58.4 59.8 48.9 72.8 60.4 60.5
InternVL3-14B zhu2025internvl3 60.3 76.0 50.9 56.2 70.3 59.3 62.2 70.7 64.0 69.8 69.3 68.5
InternVL3.5-14B 57.6 73.7 51.3 58.0 68.0 59.0 61.3 70.5 63.2 73.0 70.9 69.4
Kimi-VL-A3B-2506 team2025kimi 56.8 64.6 63.9 72.4 54.5 64.8 69.3 65.3
InternVL3.5-20B-A4B 59.0 74.2 50.5 58.5 66.6 54.4 60.5 71.2 60.0 69.6 71.5 68.1
InternVL3.5-30B-A3B 60.4 72.8 55.1 53.1 66.6 59.0 61.2 72.3 64.8 75.8 70.7 70.9
Gemma3-27B team2025gemma 54.4 55.2 61.8 65.1 51.6 79.8 61.6 64.5
Qwen2.5-VL-32B bai2025qwen2_5 61.2 64.5 66.4 71.2 60.3 85.2 69.5 71.6
Cambrian-34B tong2024cambrian 67.8 44.1
InternVL3-38B zhu2025internvl3 64.0 77.9 57.4 63.8 71.8 62.3 66.2 75.6 67.3 71.6 73.3 72.0
InternVL3.5-38B 60.9 77.4 58.9 63.7 71.8 71.9 67.4 75.9 66.0 80.0 73.1 73.8
GPT-4V gpt4v 54.6 62.7 62.3 64.3 53.1 61.4 71.8 65.6
GPT-4o-20240513 chatgpt4o 68.0 55.7 68.0 65.4 75.4 45.2 80.6 77.7 69.7
Claude-3.5-Sonnet claude3series2024 53.4 60.1 51.6
Gemini-1.5-Pro reid2024gemini1_5 53.4 64.5 67.5 38.2
GLM-4.5V hong2025glm_v_thinking 65.3 75.3 70.9 75.2 61.7 79.0 72.9 72.2
Qwen2-VL-72B wang2024qwen2vl 71.8 77.8
Qwen2.5-VL-72B bai2025qwen2_5 64.4 70.7 75.7 63.2
InternVL3-78B zhu2025internvl3 66.3 79.3 60.4 64.5 73.2 64.3 68.0 78.0 65.4 73.6 77.4 73.6
InternVL3.5-241B-A28B 61.4 77.0 61.3 47.5 72.7 73.0 65.5 75.2 65.1 82.8 75.4 74.6
Table 5: Comparison of multi-image and real-world understanding performance. Multi-image benchmarks include BLINK fu2024blink , Mantis-Eval jiang2024mantis , MMIU meng2024mmiu , MuirBench wang2024muirbench , MMT-Bench mmtbench , and MIRB zhao2024mirb . Real-world benchmarks include RealWorldQA realworldqa , MME-RealWorld zhang2024mme , WildVision lu2024wildvision , and R-Bench li2024r . Part of the results are sourced from the benchmark papers and the OpenCompass leaderboard opencompass2023 .
Model MME (sum) MMB v1.1 (EN) MMVet (turbo) MMStar Overall HallBench (avg) CRPE (relation) POPE (avg) Overall
InternVL3-1B zhu2025internvl3 1934.4 69.9 59.5 51.5 62.5 41.4 64.0 90.7 65.4
InternVL3.5-1B 1910.2 69.9 56.5 51.9 61.6 41.0 68.4 86.8 65.4
Qwen2-VL-2B wang2024qwen2vl 1872.0 72.2 49.5 48.0 59.1 41.7
Qwen2.5-VL-3B bai2025qwen2_5 2157.0 46.3 73.6
InternVL3-2B zhu2025internvl3 2221.2 78.6 62.2 60.7 70.2 42.5 71.5 89.6 67.9
InternVL3.5-2B 2123.3 76.6 71.7 62.7 71.7 48.6 75.6 87.2 70.5
MiniCPM-V-4-4B yao2024minicpm 2167.7 79.1 56.6 59.0 68.0 46.9 74.3 82.4 67.9
InternVL3.5-4B 2272.3 80.3 76.6 65.0 75.8 44.8 75.0 88.9 69.6
Qwen2-VL-7B wang2024qwen2vl 2326.8 80.7 62.0 60.7 71.6 50.6 74.4 88.1 71.0
Qwen2.5-VL-7B bai2025qwen2_5 2347.0 82.6 67.1 63.9 74.4 52.9 76.4 86.4 71.9
MiniCPM-V2.6 yao2024minicpm 2348.4 78.0 60.0 57.5 69.8 48.1 75.2 87.3 70.2
Keye-VL-8B team2025keye_vl 2214.7 76.3 70.0 72.8 74.5 57.7 76.5 87.0 73.7
GLM-4.1V-9B hong2025glm_v_thinking 2445.8 85.8 66.4 72.9 78.1 63.2 78.9 87.1 76.4
InternVL3-8B zhu2025internvl3 2415.4 81.7 81.3 68.2 79.4 49.9 76.3 91.1 72.4
InternVL3-9B zhu2025internvl3 2372.8 81.7 76.2 66.3 77.2 51.2 75.0 90.4 72.2
InternVL3.5-8B 2380.6 79.5 83.1 69.3 79.2 54.5 75.1 88.7 72.8
Gemma3-12B team2025gemma 2044.7 71.8 64.9 56.1 66.5 47.2 69.1 85.2 67.2
InternVL3-14B zhu2025internvl3 2478.3 83.5 80.2 68.8 80.3 55.1 77.3 90.2 74.2
InternVL3.5-14B 2398.4 81.5 81.7 70.4 79.8 54.0 76.6 87.7 72.8
Kimi-VL-A3B-2506 team2025kimi 2352.7 84.4 78.1 70.4 79.2 59.8 60.9 86.5 69.1
InternVL3.5-20B-A4B 2318.1 85.2 80.3 70.4 79.7 52.0 76.8 89.4 72.7
InternVL3.5-30B-A3B 2461.9 84.8 85.5 72.0 82.6 53.8 77.6 89.6 73.7
Gemma3-27B team2025gemma 1816.9 77.2 68.3 61.7 68.0 47.9 71.64 85.8 68.4
Qwen2.5-VL-32B bai2025qwen2_5 2402.9 85.2 69.6 67.8 77.1 53.7 77.0 86.8 72.5
Cambrian-34B tong2024cambrian 78.3 53.2 54.2 41.6
InternVL3-38B zhu2025internvl3 2523.6 86.9 83.9 71.5 83.1 57.1 77.1 90.6 74.9
InternVL3.5-38B 2492.4 87.3 82.2 75.3 83.5 59.7 77.7 90.4 75.9
GPT-4V gpt4v 1926.6 80.0 67.5 56.0 68.1 46.5
GPT-4o-20240513 chatgpt4o 83.1 69.1 64.7 55.0 76.6 86.9 72.8
Claude-3-Opus claude3series2024 1586.8 60.1 51.7 45.7 53.5 37.8
Claude-3.5-Sonnet claude3series2024 80.9 70.1 65.1 55.5
Gemini-1.5-Pro reid2024gemini1_5 74.6 64.0 59.1 45.6
GLM-4.5V hong2025glm_v_thinking 2423.8 88.2 75.2 75.3 81.3 64.5 55.4 85.9 68.6
Qwen2-VL-72B wang2024qwen2vl 2482.7 85.9 74.0 68.3 79.2 58.1
Qwen2.5-VL-72B bai2025qwen2_5 2448.0 88.4 76.2 70.8 80.7 55.2 79.2
InternVL3-78B zhu2025internvl3 2549.8 87.7 81.3 72.5 83.1 59.1 79.2 90.3 76.2
InternVL3.5-241B-A28B 2525.9 87.4 81.2 77.9 84.2 57.3 78.0 90.7 75.3
Table 6: Comparison of comprehensive multimodal understanding and hallucination performance. Comprehensive multimodal benchmarks include MME fu2023mme , MMBench liu2023mmbench , MMVet yu2023mmvet , and MMStar chen2024mmstar . Hallucination-related benchmarks encompass HallusionBench guan2023hallusionbench , CRPE wang2024allseeingv2 , and POPE li2023pope . Part of the results are sourced from the benchmark papers and the OpenCompass leaderboard opencompass2023 . When calculating Overall, the score of MME is normalized from 0-2800 to 0-100.
Model Name RefCOCO RefCOCO+ RefCOCOg Overall
val test-A test-B val test-A test-B val test
Grounding-DINO-L (grounding_dino, ) 90.6 93.2 88.2 82.8 89.0 75.9 86.1 87.0 86.6
UNINEXT-H (uninext, ) 92.6 94.3 91.5 85.2 89.6 79.8 88.7 89.4 88.9
ONE-PEACE (one-peace, ) 92.6 94.2 89.3 88.8 92.2 83.2 89.2 89.3 89.8
InternVL3-1B zhu2025internvl3 85.8 90.1 81.7 76.6 84.1 69.2 82.8 82.6 81.6
InternVL3.5-1B 85.4 89.7 80.2 77.7 85.5 69.5 81.9 81.6 81.4
InternVL3-2B zhu2025internvl3 89.8 92.6 86.4 84.0 89.2 76.5 87.6 87.2 86.7
InternVL3.5-2B 88.7 91.6 84.8 82.7 88.4 76.6 85.6 85.5 85.5
Qwen2.5-VL-3B bai2025qwen2_5 89.1 91.7 84.0 82.4 88.0 74.1 85.2 85.7 85.0
InternVL3.5-4B 92.5 94.3 88.2 87.6 92.3 81.6 89.6 89.3 89.4
Shikra-7B (chen2023shikra, ) 87.0 90.6 80.2 81.6 87.4 72.1 82.3 82.2 82.9
CogVLM-Grounding (wang2023cogvlm, ) 92.8 94.8 89.0 88.7 92.9 83.4 89.8 90.8 90.3
Qwen2-VL-7B wang2024qwen2vl 91.7 93.6 87.3 85.8 90.5 79.5 87.3 87.8 87.9
Qwen2.5-VL-7B bai2025qwen2_5 90.0 92.5 85.4 84.2 89.1 76.9 87.2 87.2 86.6
TextHawk2 (yu2024texthawk2, ) 91.9 93.0 87.6 86.2 90.0 80.4 88.2 88.1 88.2
InternVL3-8B zhu2025internvl3 92.5 94.6 88.0 88.2 92.5 81.8 89.6 90.0 89.6
InternVL3-9B zhu2025internvl3 91.8 93.2 86.6 86.4 91.0 79.9 88.0 88.5 88.2
InternVL3.5-8B 92.4 94.7 88.7 87.9 92.4 82.4 89.6 89.4 89.7
Ferret-v2-13B (ferretv2, ) 92.6 95.0 88.9 87.4 92.1 81.4 89.4 90.0 89.6
InternVL3-14B zhu2025internvl3 92.0 94.4 87.8 87.4 92.1 81.5 88.6 89.3 89.1
InternVL3.5-14B 92.6 94.7 89.4 88.3 92.7 82.5 90.1 90.5 90.1
InternVL3.5-20B-A4B 91.9 94.1 88.8 87.6 92.0 82.7 89.1 90.0 89.5
InternVL3.5-30B-A3B 93.1 95.4 90.1 89.6 93.2 84.4 90.6 91.0 90.9
InternVL3-38B zhu2025internvl3 93.2 95.1 90.2 89.8 93.2 85.2 91.4 91.5 91.2
InternVL3.5-38B 90.3 91.8 89.0 87.5 90.0 84.7 89.7 89.9 89.1
Qwen2-VL-72B wang2024qwen2vl 93.2 95.3 90.7 90.1 93.8 85.6 89.9 90.4 91.1
Qwen2.5-VL-72B bai2025qwen2_5 92.7 94.6 89.7 88.9 92.2 83.7 89.9 90.3 90.3
InternVL3-78B zhu2025internvl3 93.4 95.4 90.3 90.1 93.8 85.3 91.5 91.5 91.4
InternVL3.5-241B-A28B 94.1 96.3 91.5 91.6 94.6 86.9 92.0 92.1 92.4
Table 7: Comparison of visual grounding performance. We evaluate InternVL3.5’s visual grounding capability on RefCOCO, RefCOCO+, and RefCOCOg datasets kazemzadeh2014referitgame ; mao2016generation . Part of the results are collected from wang2024qwen2vl .
Model MMMB Multilingual MMBench MTVQA Overall
en zh pt ar tr ru en zh pt ar tr ru (avg)
InternVL3-1B zhu2025internvl3 79.4 70.1 62.3 58.0 47.6 61.9 72.6 66.2 62.3 48.0 39.5 60.3 22.2 47.9
InternVL3.5-1B 77.0 73.1 67.2 59.0 53.5 66.3 71.2 66.3 61.7 45.8 45.7 60.2 22.9 49.1
Qwen2-VL-2B wang2024qwen2vl 78.3 74.2 72.6 68.3 61.8 72.8 72.1 71.1 69.9 61.1 54.4 69.3 20.0 52.6
Qwen2.5-VL-3B bai2025qwen2_5 24.8
InternVL3-2B zhu2025internvl3 81.9 78.3 75.4 68.6 62.9 74.6 81.3 77.8 75.9 66.4 59.5 70.7 26.7 57.4
InternVL3.5-2B 80.2 77.7 75.9 68.5 69.1 76.3 78.4 75.9 73.7 63.7 62.0 71.4 28.5 58.0
MiniCPM-V-4-4B yao2024minicpm 82.0 80.2 75.6 60.1 63.8 71.7 80.8 80.4 71.6 51.9 59.1 67.7 22.6 54.5
InternVL3.5-4B 84.3 82.6 81.0 76.4 75.2 81.4 81.5 81.1 76.7 71.0 72.4 75.7 29.6 62.1
mPLUG-Owl2 ye2023mplug2 67.3 61.0 59.7 45.8 45.4 62.6 66.2 59.4 58.2 37.9 47.7 60.4
Qwen2-VL-7B wang2024qwen2vl 83.9 82.4 81.2 79.0 74.7 82.4 81.8 81.6 79.1 75.6 74.5 79.3 25.6 61.6
Qwen2.5-VL-7B bai2025qwen2_5 29.2
Keye-VL-8B team2025keye_vl 66.8 83.0 74.1 73.8 72.0 76.8 53.9 87.8 57.4 67.2 67.14 68.7 22.3 54.6
GLM-4.1V-9B hong2025glm_v_thinking 82.6 83.6 79.4 80.4 80.4 82.9 83.0 86.0 79.8 78.8 78.5 82.0 25.5 62.8
InternVL3-8B zhu2025internvl3 85.1 83.1 82.5 81.6 76.2 83.4 85.5 85.6 83.2 79.2 75.9 82.6 30.2 64.7
InternVL3-9B zhu2025internvl3 84.8 83.7 80.6 69.9 68.5 80.8 86.5 85.2 79.1 64.3 68.3 79.1 27.1 60.7
InternVL3.5-8B 84.9 83.0 81.4 79.6 77.4 82.1 82.5 80.7 79.0 75.9 74.8 77.6 35.2 65.0
Gemma3-12B team2025gemma 77.6 77.3 77.3 73.4 73.9 75.9 72.1 72.5 68.3 53.8 60.5 60.4 24.4 55.0
InternVL3-14B zhu2025internvl3 85.7 84.7 83.1 83.7 79.3 83.6 86.7 85.8 83.2 81.1 80.7 83.8 31.6 66.2
InternVL3.5-14B 85.1 84.1 82.7 80.3 79.4 83.5 84.0 83.7 80.0 77.8 77.0 77.0 34.2 65.5
Kimi-VL-A3B-2506 team2025kimi 83.1 78.9 76.9 71.3 71.0 76.3 82.3 79.9 76.9 62.8 66.7 73.8 27.2 59.1
InternVL3.5-20B-A4B 85.1 83.1 83.2 82.2 80.4 83.8 85.3 84.2 82.1 78.9 79.5 82.5 28.2 64.4
InternVL3.5-30B-A3B 86.4 85.7 83.4 83.3 81.7 85.0 86.1 86.3 83.1 82.3 81.5 83.5 33.7 67.3
Gemma3-27B team2025gemma 78.9 77.8 77.8 75.4 76.1 76.9 79.0 77.2 74.5 71.8 74.0 74.1 27.5 59.9
Qwen2.5-VL-32B bai2025qwen2_5 85.2 83.0 81.9 81.8 79.9 83.5 88.1 85.9 80.2 81.6 79.7 84.5 31.4 65.7
InternVL3-38B zhu2025internvl3 86.7 85.6 84.5 84.8 82.6 85.1 89.0 89.3 87.1 84.6 84.3 87.4 32.4 68.1
InternVL3.5-38B 86.7 85.5 85.1 84.1 84.3 85.3 87.4 86.9 84.2 82.0 83.4 85.6 36.1 68.7
GPT-4V gpt4v 75.0 74.2 71.5 73.5 69.0 73.1 77.6 74.4 72.5 72.3 70.5 74.8 22.0 56.1
GPT-4o chatgpt4o 27.8
Gemini-1.0-Pro team2023gemini 75.0 71.9 70.6 69.9 69.6 72.7 73.6 72.1 70.3 61.1 69.8 70.5
GLM-4.5V hong2025glm_v_thinking 87.1 86.9 84.8 84.5 84.6 84.3 89.1 89.3 86.9 83.7 84.0 87.2 30.5 67.5
Qwen2-VL-72B wang2024qwen2vl 86.8 85.3 85.2 84.8 84.2 85.3 86.9 87.2 85.8 83.5 84.4 85.3 30.9 67.2
Qwen2.5-VL-72B bai2025qwen2_5 31.7
InternVL3-78B zhu2025internvl3 87.2 86.6 85.5 86.5 84.6 86.1 89.4 90.3 88.7 86.1 86.6 88.1 32.5 68.9
InternVL3.5-241B-A28B 87.6 86.4 85.3 84.2 85.1 86.0 88.9 87.7 87.0 86.5 86.7 87.6 39.3 70.8
Table 8: Comparison of multimodal multilingual performance. We evaluate multilingual capabilities across 3 benchmarks, including MMMB sun2024parrot , Multilingual MMBench sun2024parrot and MTVQA tang2024mtvqa with six languages: English (en), Chinese (zh), Portuguese (pt), Arabic (ar), Turkish (tr), and Russian (ru).
Model Name Video-MME (wo / w sub) MVBench MMBench-Video (val) MLVU (M-Avg) LongVideoBench (val total) Overall
InternVL3-1B zhu2025internvl3 51.0 / 53.0 63.1 1.30 53.0 48.1 51.9
InternVL3.5-1B 52.4 / 55.0 61.0 1.39 56.6 53.0 54.1
Qwen2-VL-2B wang2024qwen2vl 55.6 / 60.4 63.2
Qwen2.5-VL-3B bai2025qwen2_5 61.5 / 67.6 67.0 1.63 68.2 43.3 60.3
InternVL3-2B zhu2025internvl3 58.9 / 61.4 70.4 1.42 64.2 55.4 59.6
InternVL3.5-2B 58.4 / 61.9 65.9 1.56 64.4 57.4 60.0
MiniCPM-V-4-4B yao2024minicpm 61.2 / 65.8 58.7
InternVL3.5-4B 65.4 / 68.6 71.2 1.59 70.4 60.8 64.9
VideoChat2-HD li2023videochat 45.3 / 55.7 62.3 1.22 47.9
LLaVA-OneVision-7B li2024llavaov 58.2 /   – 56.7
MiniCPM-V-2.6 yao2024minicpm 60.9 / 63.6 1.70 54.9
Qwen2-VL-7B wang2024qwen2vl 63.3 / 69.0 67.0 1.44 55.6
Qwen2.5-VL-7B bai2025qwen2_5 65.1 / 71.6 69.6 1.79 70.2 45.3 63.6
Keye-VL-8B team2025keye_vl 67.7 /   – 64.8
GLM-4.1V-9B team2025keye_vl 68.2 / 73.6 68.4 1.63 71.5 65.7 67.0
InternVL3-8B zhu2025internvl3 66.3 / 68.9 75.4 1.69 71.4 58.8 66.2
InternVL3-9B zhu2025internvl3 66.7 / 68.9 74.3 1.69 70.8 62.5 66.6
InternVL3.5-8B 66.0 / 68.6 72.1 1.67 70.2 62.1 65.8
InternVL3-14B zhu2025internvl3 70.4 / 73.0 76.6 1.73 73.3 63.9 69.1
InternVL3.5-14B 67.9 / 71.0 72.8 1.73 72.1 62.7 67.4
Kimi-VL-A3B-2506 team2025kimi 67.8 / 72.6 59.7 74.2 64.5
InternVL3.5-20B-A4B 62.4 / 64.9 73.3 1.54 65.6 58.3 62.6
InternVL3.5-30B-A3B 68.7 / 71.8 72.1 1.69 73.0 63.8 67.6
Oryx-1.5-32B liu2024oryx 67.3 / 74.9 70.1 1.52 72.3
Qwen2.5-VL-32B bai2025qwen2_5 70.5 / 77.9 1.93
VILA-1.5-40B lin2024vila 60.1 / 61.1 1.61 56.7
InternVL3-38B zhu2025internvl3 72.7 / 75.0 76.9 1.81 77.8 67.3 71.7
InternVL3.5-38B 70.9 / 74.2 75.0 1.90 77.0 65.7 71.0
GPT-4V/4T openai2023gpt4 59.9 / 63.3 43.7 1.53 49.2 59.1 54.4
GPT-4o-20240513 gpt4v 71.9 / 77.2 1.63 64.6 66.7
GPT-4o-20240806 chatgpt4o 1.87
Gemini-1.5-Pro reid2024gemini1_5 75.0 / 81.3 1.30 64.0
GLM-4.5V li2024llavaov 74.6 / 80.7 73.0 2.05 75.3 68.8 73.5
VideoLLaMA2-72B cheng2024videollama2 61.4 / 63.1 62.0
LLaVA-OneVision-72B li2024llavaov 66.2 / 69.5 59.4 66.4 61.3
Qwen2-VL-72B wang2024qwen2vl 71.2 / 77.8 73.6 1.70
Qwen2.5-VL-72B bai2025qwen2_5 73.3 / 79.1 70.4 2.02 74.6 60.7 70.9
InternVL3-78B zhu2025internvl3 72.7 / 75.7 78.7 1.81 79.5 65.7 72.1
InternVL3.5-241B-A28B 72.9 / 76.0 76.5 1.74 78.2 67.1 71.4
Table 9: Comparison of video understanding performance. We evaluate InternVL3.5’s video understanding capabilities across 5 benchmarks. For Video-MME fu2024video , MMBench-Video fang2024mmbench , MLVU MLVU , and LongVideoBench wu2024longvideobench , we test with four different settings: 16, 32, 48, and 64 frames, and report the maximum results. For MVBench li2024mvbench , we conduct testing using 16 frames. When calculating Overall, the score of MMBench-Video is normalized from 0-3 to 0-100.
3.4 Multi-Image Understanding
To assess InternVL3’s ability to understand and reason over multiple images —– a key aspect of multimodal interaction —– we conduct comprehensive evaluations on a suite of widely recognized benchmarks, including BLINK fu2024blink , Mantis-Eval jiang2024mantis , MMIU meng2024mmiu , MuirBench wang2024muirbench , MMT-Bench mmtbench , and MIRB zhao2024mirb . These benchmarks evaluate critical skills such as cross-image reasoning and context integration, which are essential for effective multimodal systems.
As shown in Table 5, across various model scales, InternVL3.5 consistently outperforms other open-source and closed-source counterparts, including earlier versions such as InternVL3. For example, InternVL3.5-38B achieves an overall score of 67.4, which is higher than the 66.2 achieved by InternVL3-38B. On the lightweight scale, InternVL3.5-2B achieves an overall score of 50.5, and its performance on individual benchmarks is 51.3 on BLINK, 58.5 on Mantis, 44.9 on MMIU, 44.0 on Muir, 58.5 on MMT and 45.9 on MIRB. Furthermore, larger model sizes lead to significant improvements in multi-image understanding capabilities. When the model is scaled up to InternVL3.5-4B, the overall score increases to 57.2, with scores of 58.1 on BLINK, 62.7 on Mantis, 49.2 on MMIU, 53.1 on Muir, 64.3 on MMT, and 55.9 on MIRB. As the model size continues to grow, performance across all benchmarks improves consistently, with InternVL3.5-8B achieving an overall score of 59.9, InternVL3.5-14B reaching 61.3, and InternVL3.5-241B-A28B improving to 65.5.
3.5 Real-World Comprehension
To evaluate the performance of InternVL3.5 on realistic and complex tasks, we provide experimental results on four real-world comprehension benchmarks: RealWorldQA realworldqa , MME-RealWorld zhang2024mme , WildVision lu2024wildvision , and R-Bench li2024r .
As shown in Table 5, InternVL3.5 achieves comparable or superior performance compared to existing methods, e.g., Qwen2.5-VL, MiniCPM-V-4 and Keye-VL. For example, the smallest variant InternVL3.5-1B demonstrates promising performance with a RealWorldQA score of 57.6, an MME-RealWorld score of 46.8, a WildVision win rate of 49.2, and an R-Bench score of 57.4. Scaling up the model results in further improvements, as larger models provide more robust representations and stronger comprehension capabilities in real-world scenarios.
At the higher end of the scale, the InternVL3.5-38B and InternVL3.5-241B-A28B models achieve top-tier results among the InternVL3.5 series. In particular, InternVL3.5-241B-A28B records an overall score of 74.6. Compared to competitive models, such as GPT-4o gpt4v —which scores 45.2 on MME-RealWorld and 80.6 on WildVision—the InternVL3.5 series exhibits competitive strengths. InternVL3.5-241B-A28B not only surpasses GPT-4o on RealWorldQA and closely matches its R-Bench performance but also considerably outperforms it on MME-RealWorld, indicating a robust overall performance on tasks demanding both perceptual precision and comprehensive understanding.
3.6 Comprehensive Multimodal Understanding
In Table 6, we evaluate InternVL3.5 on a set of comprehensive multimodal understanding benchmarks, including MME fu2023mme , MMBench (English and Chinese) liu2023mmbench , MMBench v1.1 (English) liu2023mmbench , and MMVet yu2023mmvet and MMStar chen2024mmstar .
We observe that InternVL3.5 outperforms existing methods like Keye-VL, Qwen2.5-VL and MiniCPM-V-4, especially on MMStar and MMVet. For instance, InternVL3.5-4B achieves an MMVet score of 76.6 and MMStar of 65.0, compared to 56.6 and 59.0 of MiniCPM-V-4. The improvements remain significant as model size grows, where InternVL3.5-241B-A28B finally achieves 87.4 on MMBench v1.1, 81.2 on MMVet, 77.9 on MMStar, and an overall score of 84.2.
We note that InternVL3.5 does not achieve a notable improvement compared to InternVL3. This is partly because the model’s understanding performance has approached saturation, and also partly stems from our optimization of text and reasoning capabilities—which, while achieving improvements on relevant benchmarks, slightly impairs the performance of multimodal understanding.
3.7 Multimodal Hallucination Evaluation
To evaluate the propensity for hallucination of InternVL3.5, we conduct experiments on three established benchmarks: HallusionBench guan2023hallusionbench , CRPE wang2024allseeingv2 , and POPE li2023pope . The results are shown in Table 6. Compared with previous InternVL series, the new InternVL3.5 models provide consistent improvements in handling multimodal hallucination challenges across various model scales, e.g., +2.6 on 2B scale and +1.0 on 38B scale on the overall score. Despite these advancements, there are minor declines on some model scales such as 14B and 241B-A28B, indicating that further enhancement on data and training strategies is needed to achieve more consistent improvements on all model scales, which remains an important future direction to build a more trustworthy multimodal model.
3.8 Visual Grounding
For the visual grounding task, we evaluate InternVL3.5 on RefCOCO, RefCOCO+, and RefCOCOg datasets kazemzadeh2014referitgame ; mao2016generation . As shown in Table 7, InternVL3.5 maintains the strong capabilities of InternVL3, which already achieves the upper bound of this tasks, i.e., approximately 90% average accuracy. However, the InternVL3.5 training scheme still provides additional gains on several model sizes. For example, InternVL3.5-14B achieves an overall score of 90.1 on the RefCOCO series, outperforming InternVL3-14B by +0.8%. In addition, InternVL3.5-241B-A28B builds a new state-of-the-art performance on RefCOCO, with an overall score of 92.4, further highlighting the potential of InternVL3.5 for real-world applications requiring precise multimodal understanding.
3.9 Multimodal Multilingual Understanding
InternVL3.5 exhibits strong multimodal multilingual understanding across a variety of benchmarks and languages. As summarized in Table 8, InternVL3.5 consistently achieves high scores on MMMB sun2024parrot , Multilingual MMBench sun2024parrot and MTVQA tang2024mtvqa , covering six languages including English, Chinese, Portuguese, Arabic, Turkish, and Russian. Compared to InternVL3, InternVL3.5 has significant improvements in its language capabilities, thus achieving better results on these multilingual benchmarks. For example, InternVL3.5-1B achieves up to 1.2% gains over InternVL3-1B. Compared to other leading multimodal models, such as Qwen2.5-VL and GPT-4V, InternVL3.5 also demonstrates notable improvements in both overall accuracy and language coverage, especially at larger scales. For example, InternVL3.5-241B-A28B outperforms GPT-4V by +14.7% on the overall score across all multilingual benchmarks. The results highlight InternVL3.5’s robust capability to handle complex multilingual and multimodal tasks, making it highly effective for global applications that require comprehensive cross-language understanding.
3.10 Video Understanding
InternVL3.5 demonstrates remarkable video understanding capabilities across a comprehensive set of benchmarks. As presented in Table 9, InternVL3.5 consistently achieves competitive or leading scores on Video-MME fu2024video , MVBench li2024mvbench , MMBench-Video fang2024mmbench , MLVU MLVU , LongVideoBench wu2024longvideobench . Performance improvements are observed across almost all metrics for small-size models. In particular, InternVL3.5-1B outperforms InternVL3-1B by +2.2% of overall performance. For larger InternVL3.5 variants (such as 38B) also deliver comparable results with other state-of-the-art models of similar scale. Furthermore, InternVL3.5 exhibits robust generalization on challenging tasks involving long video sequences and complex reasoning, as reflected in its performance on LongVideoBench. In particular, InternVL3.5-1B achieves significant improvements on LongVideoBench, i.e., +4.9%. The model’s ability to process multi-frame inputs and handle diverse video scenarios underscores its versatility. These results highlight InternVL3.5’s substantial progress in video understanding, positioning it as a highly capable solution for advanced multimodal video analysis tasks.
Model ScreenSpot ScreenSpot-v2 OSWorld-G WindowsAgentArena WebArena-Lite-v2
ShowUI-2B lin2024showui 75.1
UI-TARS-2B qin2025ui 82.3 84.7
JEDI-3B xie2025scalingcomputerusegroundinguser 80.9 27.3
OS-Atlas-4B wu2024atlas 70.1 71.9
Qwen2.5-VL-3B bai2025qwen2_5 80.9 27.3
InternVL3.5-4B 83.6 85.1 33.9 9.7 7.8
OS-Atlas-7B wu2024atlas 82.5 84.1 47.5
UGround-V1-7B gou2025uground 86.3 36.4
Aguvis-7B xu2024aguvis 81.8 38.7
UI-TARS-7B qin2025ui 89.5 91.6 47.5
UI-TARS-1.5-7B qin2025ui 89.7 64.2 15.9 17.5
JEDI-7B xie2025scalingcomputerusegroundinguser 91.7 54.1
Qwen2.5-VL-7B bai2025qwen2_5 88.8 31.4 3.4
GLM-4.1V-9B-Thinking† (hong2025glm_v_thinking, )
MiMo-VL-7B-RL coreteam2025mimovltechnicalreport 87.2 90.5 50.7
InternVL3-8B zhu2025internvl3 79.5 81.4
InternVL3.5-8B 87.9 86.2 36.4 10.5 12.3
InternVL3.5-14B 87.5 88.6 44.7 12.5 12.3
GTA1-32B yang2025gta1 93.2 61.9
Qwen2.5-VL-32B bai2025qwen2_5 91.3 46.5
Gemma-3-27B-IT† team2025gemma
InternVL3-38B zhu2025internvl3 85.6 88.3
InternVL3.5-20B-A4B 85.5 87.6 38.2 11.0 9.5
InternVL3.5-30B-A3B 86.6 87.3 42.4 11.0 10.4
InternVL3.5-38B 81.0 83.5 42.9 14.5 7.1
Operator‡ cua2025 70.5 40.6
Aguvis-72B xu2024aguvis 89.2 3.5 9.0
UI-TARS-72B qin2025ui 88.4 90.3 57.1 17.9 10.3
GPT-4o chatgpt4o 18.1 3.5 1.9
Claude-3.5-Sonnet claude3series2024 83.0
Claude-3.7-Sonnet anthropic2025claude37 87.6 6.4 1.9
Gemini-2.0-Flash gemini2_0 84.0
Gemini-2.5-Pro geminipro2.5 45.2
Seed1.5-VL guo2025seedvl1_5 95.2 62.9
Qwen2.5-VL-72B bai2025qwen2_5 87.1 9.7 14.4
InternVL3-78B zhu2025internvl3 88.7 90.9
InternVL3.5-241B-A28B 89.8 92.9 53.2 18.0 11.7
Table 10: Comparison of GUI grounding and online agentic evaluation results. To assess the GUI agent capabilities of InternVL3.5, we conducted evaluations on a diverse set of platforms. We evaluate GUI grounding capabilities across 3 benchmarks including ScreenSpot cheng2024seeclick , ScreenSpot-v2 wu2024atlas and OSWorld-G xie2025scalingcomputerusegroundinguser . For online agentic evaluation, our assessment covers Ubuntu, Windows, and Web utilizing the OSWorld xie2024osworld , WindowsAgentArena bonatti2024windows , and WebArena-Lite-v2 wang2025mmbenchgui . Models with symbols † and ‡  are evaluated under 100 and 200 steps, respectively, while all other results were evaluated under 50 steps.
3.11 GUI Agent Tasks
To validate the GUI agent capabilities of InternVL3.5, we conduct detailed experiments in Table 10. In particular, we evaluate InternVL3.5 on six GUI grounding and agent tasks, namely ScreenSpot cheng2024seeclick , ScreenSpot-v2 wu2024atlas , OSWorld-G xie2024osworld , OSWorld xie2024osworld , WindowsAgentArena bonatti2024windows , and WebArena-Lite-v2 wang2025mmbenchgui . In GUI grounding tasks, InternVL3.5 outperforms most open-source models and is close to the performance of closed-source models. For example, InternVL3.5-241B-A28B outperforms the specialized model, e.g., +2.6% against UI-TARS-72B qin2025ui on ScreenSpot-v2. Compared to the most advanced commercial model, i.e., Seed1.5-VL guo2025seedvl1_5 , InternVL3.5-241B-A28B still maintains close performance on ScreenSpot-v2, i.e., 92.9 vs. 95.2. In GUI agent tasks, InternVL also demonstrates competitive results on different platforms. In particular, InternVL3.5-241B-A28B achieves the best results against existing generalist MLLMs, e.g., +8.3% over Qwen2.5-VL-72B on WindowsAgentArena. In particular, GPT-4o only achieves a score of 3.5 on this challenging benchmark. On WebArena-Lite-v2, the performance score of GPT-4o further decreases to 1.9, but InternVL3.5-241B-A28B can still achieve a top-tier score of 11.7. These results confirm the great potential of InternVL3.5 as a fundamental model for GUI tasks.
Method VSI-Bench ERQA SpaCE-10 OmniSpatial Overall
InternVL3-1B zhu2025internvl3 29.7 30.3 41.3 35.5 34.2
InternVL3.5-1B 49.3 35.3 33.6 40.7 39.7
Qwen2.5-VL-3B bai2025qwen2_5 27.9 38.0 34.8 40.3 35.3
InternVL3-2B zhu2025internvl3 31.5 31.5 44.2 38.0 36.3
InternVL3.5-2B 53.8 37.3 34.6 42.3 42.0
MiniCPM-V-4-4B yao2024minicpm 30.3 36.3 39.0 43.1 37.2
InternVL3.5-4B 54.9 38.5 35.5 45.8 43.7
Qwen2.5-VL-7B bai2025qwen2_5 35.9 38.8 33.3 39.2 36.8
MiMo-VL-RL-8B coreteam2025mimovltechnicalreport 36.4 37.8 36.1 46.5 39.2
Keye-VL-8B team2025keye_vl 28.6 35.3 38.6 46.5 37.2
GLM-4.1V-9B hong2025glm_v_thinking 39.2 45.8 43.4 47.7 44.0
InternVL3-8B zhu2025internvl3 42.1 35.3 40.0 41.6 39.7
InternVL3.5-8B 56.3 41.0 39.5 47.8 46.1
Gemma-3-12B team2025gemma 21.9 36.1 41.5 43.7 35.8
InternVL3-14B zhu2025internvl3 48.9 39.5 47.3 45.9 45.4
InternVL3.5-14B 60.8 41.8 48.8 47.6 49.7
Kimi-VL-A3B-2506 team2025kimi 37.4 36.0 39.2 37.3 37.5
InternVL3.5-20B-A4B 60.1 41.6 51.6 45.4 46.7
InternVL3.5-30B-A3B 63.7 41.5 49.7 48.1 50.8
Gemma-3-27B team2025gemma 22.0 37.5 41.5 44.8 36.4
Qwen2.5-VL-32B bai2025qwen2_5 34.7 40.7 32.6 47.4 38.8
InternVL3-38B zhu2025internvl3 48.9 42.8 53.1 48.5 48.3
InternVL3.5-38B 66.3 43.3 43.8 51.4 51.2
GPT-4o-20241120 chatgpt4o 34.0 47.0 49.0 47.8 44.5
GPT-5-20250807 gpt5 37.5 65.7 43.8 59.6 51.7
Gemini-2.5-Pro gemini2_0 47.8 48.3 52.7 55.2 51.0
Claude-3.7-Sonnet claude3series2024 47.0 35.5 46.2 46.9 43.9
GLM-4.5V hong2025glm_v_thinking 41.4 46.5 51.6 51.0 47.6
Qwen2.5-VL-72B bai2025qwen2_5 36.1 44.8 37.9 47.9 41.7
Step3-321B-A38B wang2025step 34.2 44.5 42.6 47.0 42.1
InternVL3-78B zhu2025internvl3 48.4 45.9 52.5 49.3 49.0
InternVL3.5-241B-A28B 69.5 46.8 55.0 51.9 55.8
Table 11: Comparison of embodied task performance. We compare InternVL3.5 and existing methods on VSI-Bench yang2024think , ERQA erqa , SpaCE-10 gong2025space10 , and OmniSpatial jia2025omnispatial . For SpaCE-10 gong2025space10 , we report the single-choice performance.
Model Semantics $\uparrow$ Count $\uparrow$ Color $\uparrow$ Shape $\uparrow$ Reasoning $\uparrow$ Overall $\uparrow$
Gemma-1.1-2B team2024gemma 32.1 33.3 25.0 35.6 28.7 31.7
InternVL3.5-1B 25.5 22.7 24.9 24.4 30.4 25.0
InternVL3.5-2B 25.2 20.1 45.4 33.4 26.5 30.7
InternVL3.5-4B 41.2 55.2 81.9 62.4 39.4 57.7
InternLM2.5-7B cai2024internlm2 27.3 31.7 59.8 51.5 28.2 42.1
Keye-VL-8B team2025keye_vl 41.4 47.5 71.4 54.9 40.6 52.2
GLM-4.1V-9B hong2025glm_v_thinking 41.6 55.6 79.1 61.5 40.0 57.1
InternVL3-8B zhu2025internvl3 33.7 46.5 69.8 59.1 36.1 50.6
InternVL3.5-8B 39.7 54.0 82.3 53.4 41.7 54.9
Gemma-3-12B team2025gemma 24.8 30.8 47.2 25.7 22.8 30.5
DeepSeek-Coder-V2-16B zhu2024deepseek_coder_v2 30.9 37.9 63.7 54.8 26.8 45.1
InternVL3-14B zhu2025internvl3 38.2 52.9 74.4 54.1 41.7 52.9
InternVL3.5-14B 44.3 55.3 77.8 63.7 45.1 58.5
Kimi-VL-A3B-2506 team2025kimi 31.1 41.5 67.0 47.4 32.4 44.9
InternVL3.5-20B-A4B 51.2 60.6 89.8 74.7 55.5 67.6
InternVL3.5-30B-A3B 51.8 66.8 91.8 75.7 53.8 69.4
Gemma-3-27B team2025gemma 36.7 51.4 76.3 62.1 39.4 54.7
Qwen2.5-VL-32B bai2025qwen2_5 40.0 55.7 76.3 61.2 43.9 56.5
InternVL3-38B zhu2025internvl3 40.8 58.7 82.2 63.6 43.9 59.1
InternVL3.5-38B 47.6 66.5 90.5 80.4 54.6 69.5
GPT-5 gpt5 67.8 72.6 91.7 81.9 68.7 77.5
GLM-4.5V hong2025glm_v_thinking 47.3 63.7 87.3 72.3 55.8 66.1
Qwen2.5-VL-72B bai2025qwen2_5 40.2 55.1 80.1 62.0 41.1 57.1
Step3-321B-A38B wang2025step 35.9 54.0 82.8 63.2 38.6 56.5
InternVL3-78B zhu2025internvl3 41.0 59.1 84.0 65.2 47.0 60.3
InternVL3.5-241B-A28B 51.2 69.2 92.1 77.6 58.0 70.7
Table 12: Comparison of SVG understanding performance on SGP-Bench qiu2024can .
Model Text2SVG Img2SVG
FID $\downarrow$ FID-C $\downarrow$ CLIP $\uparrow$ DINO $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$
InternVL3.5-1B 22.50 12.16 72.43 0.79 0.57 0.35 7.35
InternVL3.5-2B 20.98 11.26 72.71 0.81 0.56 0.34 7.44
InternVL3.5-4B 17.06 7.54 74.35 0.84 0.61 0.30 8.37
Llama-3.1-8B dubey2024llama3 19.43 11.25 71.86
Qwen2.5-VL-7B bai2025qwen2_5 24.78 15.45 71.38 0.78 0.51 0.38 6.53
Keye-VL-8B team2025keye_vl 21.96 14.39 71.17 0.80 0.53 0.37 6.94
GLM-4.1V-9B hong2025glm_v_thinking 22.68 10.45 73.20 0.82 0.54 0.35 7.33
InternVL3-8B zhu2025internvl3 23.06 14.30 71.45 0.81 0.56 0.36 7.22
InternVL3.5-8B 17.36 7.13 75.01 0.85 0.62 0.29 8.74
Llama-3.2-11B-Vision dubey2024llama3 28.16 14.35 71.49 0.76 0.47 0.39 5.91
Gemma-3-12B team2025gemma 17.14 10.41 71.62 0.82 0.58 0.35 7.63
InternVL3-14B zhu2025internvl3 19.00 13.22 71.49 0.83 0.56 0.36 7.34
InternVL3.5-14B 15.90 5.99 75.91 0.86 0.61 0.31 8.46
Kimi-VL-A3B team2025kimi 30.81 16.99 70.54 0.80 0.56 0.36 7.18
InternVL3.5-20B-A4B 16.78 5.60 77.46 0.91 0.71 0.20 12.75
InternVL3.5-30B-A3B 16.31 5.84 76.40 0.88 0.65 0.27 9.64
Gemma-3-27B team2025gemma 15.15 9.30 73.28 0.83 0.60 0.35 7.83
Qwen2.5-VL-32B bai2025qwen2_5 20.04 10.39 73.23 0.84 0.56 0.36 7.50
InternVL3-38B zhu2025internvl3 18.01 11.04 73.08 0.83 0.55 0.35 7.31
InternVL3.5-38B 14.56 5.22 76.49 0.86 0.61 0.32 8.39
Grok-3 grok3 21.97 8.69 76.80
Llama-3.1-70B dubey2024llama3 18.03 8.30 73.88
Llama-3.1-405B dubey2024llama3 16.79 8.39 73.92
DeepSeek-V3-671B-A37B liu2024deepseekv3 24.99 8.80 76.47
GPT-4o chatgpt4o 15.18 6.76 77.74 0.87 0.62 0.32 8.44
GLM-4.5V hong2025glm_v_thinking 16.64 5.09 78.35 0.87 0.63 0.32 8.67
Claude-3.7-Sonnet claude3series2024 14.38 3.50 80.79 0.91 0.65 0.29 9.26
Claude-4-Sonnet anthropic2025claude4 15.84 4.29 80.58 0.92 0.67 0.28 9.86
Gemini-2.5-Flash gemini2_5 16.72 5.21 78.22 0.88 0.59 0.32 8.32
Llama-3.2-90B-Vision dubey2024llama3 19.31 8.55 74.00 0.76 0.44 0.38 5.78
Llama-4-Scout meta2025llama4scout 17.91 9.38 73.56 0.84 0.58 0.35 7.74
Llama-4-Maverick meta2025llama4maverick 14.93 6.53 75.82 0.86 0.60 0.33 8.03
Step3-321B-A38B wang2025step 20.06 9.71 74.18 0.83 0.56 0.34 7.52
Qwen2.5-VL-72B bai2025qwen2_5 15.95 9.88 73.68 0.84 0.58 0.35 7.83
InternVL3-78B zhu2025internvl3 17.58 10.60 73.12 0.85 0.58 0.34 7.80
InternVL3.5-241B-A28B 11.27 4.43 76.81 0.88 0.64 0.29 9.19
Table 13: Comparison of SVG generation performance on SArena-Icon (Text2SVG and Img2SVG).
Dataset Version
Qwen3-0.6B
InternVL3.5-1B
Qwen3-1.7B
InternVL3.5-2B
Qwen3-4B
InternVL3.5-4B
Qwen3-8B
InternVL3.5-8B
Qwen3-14B
InternVL3.5-14B
Qwen3-30B-A3B
InternVL3.5-30B-A3B
Qwen3-32B
InternVL3.5-38B
Qwen3-235B-A22B
InternVL3.5-241B-A28B
MMLU 4d595a 52.8 49.1 62.6 61.9 73.0 73.8 76.9 78.2 81.1 81.5 81.4 83.0 83.6 84.6 87.8 89.1
CMMLU c13365 43.4 46.6 59.8 59.4 71.8 70.6 76.6 76.2 81.6 79.8 83.0 82.5 84.5 84.4 87.4 90.2
C-Eval 2daf24 42.6 49.0 61.0 61.0 72.2 71.9 77.9 77.9 81.0 80.3 82.9 83.2 83.3 85.0 86.1 90.9
GAOKAO 4c31db 40.4 48.4 64.1 68.1 80.0 83.2 85.6 84.1 90.0 87.2 89.5 92.6 93.2 93.4 95.0 94.5
TriviaQA 2121ce 18.7 20.4 30.6 31.7 39.9 40.2 52.0 49.7 60.9 55.3 58.4 57.6 63.4 59.9 73.7 74.8
NaturalQuestions 3dcea1 11.2 15.0 21.4 23.5 29.3 29.0 36.5 32.6 42.2 36.7 42.5 36.7 45.8 40.1 54.8 50.4
C3 8c358f 64.1 65.5 54.2 79.5 79.8 89.5 83.0 91.6 84.3 94.6 90.8 95.1 93.5 95.6 96.4 97.8
RACE-High 69ee4f 45.3 67.2 65.6 78.8 82.7 86.4 86.2 88.4 87.2 90.7 78.7 92.3 85.6 92.1 90.5 94.2
WinoGrande b36770 51.5 54.9 53.5 59.3 64.8 69.1 70.8 75.2 76.6 80.5 73.0 86.5 74.8 83.4 84.8 91.7
HellaSwag e42710 42.7 49.5 59.0 74.4 78.1 88.5 84.6 91.5 88.2 94.1 89.0 96.3 90.5 95.9 91.1 97.0
BBH 5b92b0 41.5 46.8 54.5 62.3 72.6 78.6 78.4 79.8 81.1 82.3 81.5 83.8 87.4 87.5 88.9 86.5
GSM8K 1d7fe4 59.6 62.8 75.4 77.2 87.8 92.0 89.8 90.9 92.5 95.6 91.8 91.4 93.4 91.5 94.4 91.6
MATH 393424 32.4 68.2 43.5 85.5 54.1 94.4 60.8 93.3 62.0 93.6 59.0 93.6 61.6 94.3 71.8 94.5
AIME2024 10.0 13.7 40.0 44.2 66.7 72.8 76.0 77.7 80.0 77.4 83.3 79.4 73.3 81.5 86.7 84.1
AIME2025 13.3 14.7 23.3 42.9 50.0 57.6 67.3 64.0 66.7 63.9 70.0 62.7 60.0 72.1 83.3 75.6
HumanEval 8e312c 39.6 45.7 72.0 65.9 82.3 87.8 85.4 93.9 89.0 97.0 89.0 96.3 89.6 98.2 91.5 98.2
Overall 38.1 44.8 52.5 61.0 67.8 74.1 74.6 77.8 77.8 80.7 77.8 82.0 79.0 83.7 85.3 87.6
Table 14: Comparison of text-related performance across multiple benchmarks. Results were obtained using the OpenCompass toolkit opencompass2023 . We compare InternVL3.5 with Qwen3 models, whose corresponding pre-trained base models are employed as the initialization of the language component in InternVL3.5. Please note that the evaluation scores of the Qwen3 series may differ from those officially reported, as we have adopted the prompt versions provided in the table across all datasets for OpenCompass evaluation.
3.12 Embodied Agent Tasks
In Table 11, we evaluate the embodied capabilities of InternVL3.5 on four benchmarks: VSI-Bench yang2024think , ERQA erqa , Space10 gong2025space10 , and OmniSpatial jia2025omnispatial . From this table, we observe the strong embodied capabilities of InternVL3.5 against previous works. On VSI-Bench, the most popular benchmark for spatial reasoning, InternVL3.5-1B achieves an overall score of 49.3, outperforming its predecessor by +19.6%. We note that InternVL3.5-1B can already achieve the state-of-the-art performance on VSI-Bench against much larger models like Qwen2.5-VL-72B bai2025qwen2_5 . When the model size of InternVL3.5 increases, the performance on VSI-Bench consistently improves from 49.3 to 69.5, greatly validating the salability of InternVL3.5 on embodied tasks. In addition to VSI-Bench, InternVL3.5 also demonstrates promising results on ERQA, Space10 and OminiSpatial. Among them, InternVL3.5B-241B-A28B scores 46.8 on ERQA, which is close to the 48.3 score of the top closed-source model Gemini-2.5-Pro. In terms of overall performance, InternVL3.5B-241B-A28B also reaches the highest score against other models, confirming its strong capabilities in embodied tasks.
3.13 SVG Tasks
Scalable vector graphics (SVG) is a common format for describing graphics on web pages, and its understanding is significant for web-based agent tasks. To evaluate this capability, We evaluate InternVL3.5 on SGP-Bench qiu2024can (Table 12), where it delivers strong results across model scales and sets new open-source state-of-the-art at larger capacities. At the small scale, InternVL3.5-4B already surpasses models such as Kimi-VL-A3B team2025kimi and even the earlier InternVL3-14B zhu2025internvl3 . In the mid-size range (8B, 14B), InternVL3.5 shows broad improvements, and the 14B variant surpasses larger counterparts such as Gemma-3-27B team2025gemma . At the high end, InternVL3.5-30B-A3B and InternVL3.5-38B achieve nearly 70% accuracy on SGP-Bench, advancing the state-of-the-art among open models and outperforming competitors such as Step3-321B-A38B wang2025step , Qwen2.5-VL-72B bai2025qwen2_5 and GLM-4.5V hong2025glm_v_thinking . Finally, InternVL3.5-241B-A28B sets a new record for open-source models, achieving the best performance across all categories except when compared to GPT-5 gpt5 .
In the SArena-Icon generation tasks (Text2SVG and Img2SVG), InternVL3.5 establishes new state-of-the-art performance among open models (Table 13). In Text2SVG, our models achieve markedly lower FID and FID-C scores than previous baselines, with the 38B variant reducing FID to 14.56. This performance even surpasses GPT-4o chatgpt4o (15.18), highlighting that InternVL3.5 is able to match or exceed the capabilities of much larger proprietary systems. In Img2SVG, the 30B and 38B variants deliver leading results on structural similarity metrics, outperforming open-source peers of comparable scale and closely matching the best proprietary systems. Furthermore, InternVL3.5-241B-A28B achieves an even stronger balance of fidelity and perceptual quality, with an FID of 11.27 and an FID-C of 4.43 in Text2SVG, together with competitive Img2SVG scores. These results highlight both the efficiency and scalability of InternVL3.5, showing that it consistently outperforms prior open models across all metrics, narrows the gap with proprietary leaders, and establishes itself as the most powerful open-source framework for SVG generation to date.
3.14 Evaluation on Language Capability
To evaluate the language capabilities of InternVL3.5, we use benchmarks covering comprehensive assessments in general knowledge (MMLU hendrycks2020measuring , CMMLU li2023cmmlu , C-Eval huang2023ceval , GAOKAO-Bench Zhang2023gaokao ), linguistic understanding (TriviaQA joshi2017triviaqa , NaturalQuestions naturalquestion , C3 sun2019investigating , RACE lai2017race ), reasoning (WinoGrande sakaguchi2020winogrande , HellaSwag zellers2019hellaswag , BigBench Hard suzgun2023bigbench ), mathematics (GSM8K-Test cobbe2021training , MATH DBLP:conf/nips/HendrycksBKABTS21 , AIME24 aime2024 , AIME25 aime2025 ), and coding (HumanEval chen2021evaluating ) tasks. As shown in Table 14, InternVL3.5 achieves even better performance than its corresponding language models on most benchmarks. Specifically, InternVL3.5-1B outperforms Qwen3-0.6B on 15 of 16 text-related benchmarks, with an overall performance gain of +6.7. For larger models such as InternVL3.5-241B-A28B, the performance improvement is also obvious, i.e., +2.3 over Qwen3-235B-A22B. These improvements come not only from the high-quality text corpora we use during pre-training and SFT, but also from our Cascade RL, which significantly benefits text-based reasoning tasks. The improvement of InternVL3.5 in text capabilities has also greatly compensated for the shortcomings of open-source multimodal models in general capabilities.
3.15 Ablation Study
Figure 5: Ablation on Cascade RL. We report average scores on the same set of multimodal reasoning and mathematical benchmarks as in Table 3. Full results are provided in Table 15.
Model MMMU (val) MathVista (mini) MathVision MathVerse (vision-only) DynaMath (worst case) WeMath LogicVista Overall
InternVL3-1B 43.4 45.8 18.8 18.7 5.8 13.4 29.8 25.1
InternVL3.5-1B-Instruct 37.2 48.6 15.8 27.0 8.4 13.9 29.1 25.7
InternVL3.5-1B-MPO 40.3 50.5 22.0 32.1 9.0 16.8 32.7 29.1
InternVL3.5-1B-CascadeRL 44.2 59.3 27.3 37.8 17.2 21.5 29.3 33.8
InternVL3-2B 48.6 57.0 21.7 25.3 14.6 22.4 36.9 32.4
InternVL3.5-2B-Instruct 53.0 60.8 27.0 39.6 19.8 28.1 41.2 38.5
InternVL3.5-2B-MPO 54.3 62.6 34.2 46.4 21.0 28.1 40.9 41.1
InternVL3.5-2B-CascadeRL 59.0 71.8 42.8 53.4 31.5 48.5 47.7 50.7
InternVL3.5-4B-Instruct 64.3 71.4 40.5 50.0 30.7 35.6 53.5 49.4
InternVL3.5-4B-MPO 65.4 71.7 48.0 54.9 30.7 39.8 55.9 52.3
InternVL3.5-4B-CascadeRL 66.6 77.1 54.4 61.7 35.7 50.1 56.4 57.4
InternVL3-8B 62.7 71.6 29.3 39.8 25.5 37.1 44.1 44.3
InternVL3.5-8B-Instruct 68.1 74.2 46.4 55.8 30.7 46.0 53.9 53.6
InternVL3.5-8B-MPO 71.2 75.9 52.6 54.8 33.1 47.7 58.6 56.3
InternVL3.5-8B-CascadeRL 73.4 78.4 56.8 61.5 37.7 57.0 57.3 60.3
InternVL3-14B 67.1 75.1 37.2 44.4 31.3 43.0 51.2 49.9
InternVL3.5-14B-Instruct 71.8 73.4 48.7 55.5 31.9 45.7 57.5 54.9
InternVL3.5-14B-MPO 73.3 74.0 53.0 57.5 32.3 45.2 60.9 56.6
InternVL3.5-14B-CascadeRL 73.3 80.5 59.9 62.8 38.7 58.7 60.2 62.0
InternVL3.5-30B-A3B-Instruct 72.3 73.3 45.1 50.4 31.9 39.7 56.4 52.7
InternVL3.5-30B-A3B-MPO 71.7 75.3 50.7 58.5 32.9 43.7 59.7 56.1
InternVL3.5-30B-A3B-CascadeRL 75.6 80.9 55.7 60.4 36.5 48.4 55.7 59.0
InternVL3-38B 70.1 75.1 34.2 48.2 35.3 48.6 58.4 52.8
InternVL3.5-38B-Instruct 73.9 75.9 58.2 59.0 29.7 47.5 60.0 57.7
InternVL3.5-38B-MPO 76.9 80.5 56.3 59.4 36.9 55.6 64.2 61.4
InternVL3.5-38B-CascadeRL 76.9 81.9 63.7 67.6 41.7 64.8 65.3 66.0
InternVL3-78B 72.2 79.0 43.1 51.0 35.1 46.1 55.9 54.6
InternVL3.5-241B-A28B-Instruct 76.2 80.1 55.6 61.7 36.5 49.7 63.3 60.4
InternVL3.5-241B-A28B-MPO 76.0 82.2 55.3 64.1 38.3 51.3 69.4 62.4
InternVL3.5-241B-A28B-CascadeRL 77.7 82.7 63.9 68.5 46.5 62.3 66.7 66.9
Table 15: Comparison of multimodal reasoning performance after different training stages.
Model GPU Hours
MMMU
Val
MathVista
MINI
MathVision
MINI
MathVerse
Vision-Only
DynaMath
WeMath LogicVista Overall
InternVL3.5-8B-Instruct 68.1 74.2 46.4 55.8 30.7 46.0 53.9 53.6
+MPO $\sim$0.3K Hours 71.2 75.9 52.6 54.8 33.1 47.7 58.6 56.3
+GSPO (1 episode) $\sim$5.5K Hours 73.8 77.9 51.6 58.8 35.1 48.9 54.8 57.3
+GSPO (2 episodes) $\sim$11.0K Hours 72.0 78.1 51.6 58.5 35.7 54.1 57.0 58.2
+CascadeRL (ours) $\sim$5.8K Hours 73.4 78.4 56.8 61.5 37.7 57.0 57.3 60.3
Table 16: Comparison of training efficiency and effectiveness of MPO, GSPO, and Cascade RL.
Cascade Reinforcement Learning (Cascade RL). To validate the effectiveness of Cascade RL, we conduct an ablation study in Figure 5 and Table 15. We evaluate a baseline model InternVL3 as well as InternVL3.5 after different training stages: InternVL3.5-Instruct (after SFT), InternVL3.5-MPO (after the first substage in Cascade RL), and InternVL3.5-CascadeRL (after both substages in Cascade RL). From it we can see that InternVL3.5-Instruct already outperforms InternVL3 by margins, e.g., + 9.3% on the 8B model. Even based on these strong SFT baselines, the performance of InternVL3.5 can be further improved after the MPO stage, providing up to +3.5% average gains on reasoning tasks. Compared to MPO-based models, our Cascade RL still provides orthogonal gains for all dense and MoE models. For example, InternVL3.5-2B obtains 12.2% average performance gains on reasoning tasks compared to the SFT model. Similar merits can also be observed on larger models, e.g., +6.5% on InternVL3.5-241B-A28B. These ablations confirm the effectiveness, stability, and scalability of our Cascade RL. Additionally, we present a comparison of the efficiency and effectiveness of our proposed Cascade RL against MPO and GSPO in Table 16. For MPO and Cascade RL, we report performance after training for one episode, whereas for GSPO, we report results after both one and two episodes. We observe that MPO achieves performance gains with only a small number of GPU hours, while GSPO yields more significant improvements but at the cost of substantially higher computational consumption. In contrast, Cascade RL attains even greater performance improvements while requiring only half the GPU hours of GSPO.
Model DocVQA ChartVQA InfoVQA TextVQA OCRBench AI2D MMStar MMMU Mathvista Overall
InternVL3.5-8B 92.3 86.7 76.2 78.2 83.2 84.0 69.3 73.4 78.4 80.2
InternVL3.5-8B-Flash 91.9 86.6 76.0 77.2 83.0 83.6 68.6 72.9 78.0 79.8
InternVL3.5-38B 94.0 88.8 81.0 82.7 87.0 87.8 75.3 76.9 81.9 83.9
InternVL3.5-38B-Flash 93.5 88.1 81.0 82.1 86.5 87.2 75.0 76.3 81.3 83.4
InternVL3.5-30B-MoE 94.2 87.4 77.8 80.5 88.0 86.8 72.0 75.6 80.9 82.6
InternVL3.5-30B-MoE-Flash 93.2 87.3 77.6 80.2 87.8 86.3 71.7 75.2 80.5 82.2
InternVL3.5-235B-MoE 94.9 88.0 82.0 84.5 90.7 86.9 77.9 77.7 82.7 85.0
InternVL3.5-235B-MoE-Flash 94.0 87.9 81.0 84.1 90.3 86.1 77.4 77.0 83.0 84.5
Table 17: Performance comparison of InternVL3.5 and InternVL3.5-Flash.
Resolution Setting Request Throughput (requests / s)
InternVL3.5-38B InternVL3.5-241B-A28B
448 Baseline 12.39 11.29
+ DvD 14.69 (1.19 $\times$) 14.05 (1.24 $\times$)
+ DvD + ViR 18.62 (1.50 $\times$) 20.84 (1.85 $\times$)
896 Baseline 2.71 2.54
+ DvD  5.06 (1.87 $\times$) 4.73 (1.86 $\times$)
+ DvD + ViR 10.97 (4.05 $\times$) 8.81 (3.47 $\times$)
1344 Baseline 1.48 1.37
+ DvD 2.92 (1.97 $\times$) 2.75 (2.01 $\times$)
+ DvD + ViR 5.14 (3.47 $\times$) 4.27 (3.12 $\times$)
Table 18: Ablation of Decoupled Vision-Language Deployment (DvD) and Visual Resolution Router (ViR) on inference efficiency. We send 16 requests per second to the deployed server. In all settings, the language models are deployed on 8 A100 GPUs.
[…]
In Table 17, we compare the performance of InternVL3.5 and InternVL3.5-Flash on several significant benchmarks. For these results, we can see that InternVL3.5-Flash can maintain the multimodal understanding and reasoning performance. In high-resolution tasks like DocVQA and InfoVQA, InternVL3.5-Flash can reach almost the 100% performance of InternVL3.5, e.g., 80.2 vs. 79.8 on 8B model size. Even when the model size improves to 241B, similar observations still hold. These results further confirm that ViR can greatly benefit the model performance without sacrificing performance.
Decoupled Vision-Language Deployment (DvD). In Table 18, we conduct detailed ablation on InternVL3.5 to show the benefit of DvD. From this table, the first observation is that DvD greatly accelerates the inference of both dense and MoE models, by up to 2.01 times and 1.97 times for InternVL3.5-241B-A28B and InternVL3.5-38B, respectively. In addition, the efficiency gains of DvD can benefit both the pre-filling and next-token prediction stages of InternVL3.5. Another finding is that as the input resolution increases, DVD efficiency also improves. For example, on InternVL3.5-38B, the speed-up of DvD can be improved from 1.19 to 1.97 as the resolution increases from 448 to 1344. This phenomenon can be attributed to the fact that larger input resolution or visual backbone leads to higher visual computational costs and further blocks the computation of the LLM. It is worth noting that the increased computational cost of high-resolution images arises because mainstream MLLMs typically adopt a dynamic high-resolution strategy, which increases the number of patches fed into the vision encoder and thereby raises the overall computation. In practical applications, beyond high-resolution images, many tasks also require multi-image and video understanding. In such scenarios, the number of patches processed by the vision encoder grows even further, leading to greater visual overhead. We believe that our proposed DvD can deliver even more significant performance gains in these scenarios.
4 Conclusion
[…]